### v1

# What Is Seen Cannot Be Unseen: The Disruptive Effect of Knowledge Conflict on Large Language Models

Abstract
Large language models frequently rely on both contextual input and parametric knowledge to perform tasks. However, these sources can come into conflict, especially when retrieved documents contradict the model’s parametric knowledge. We propose a diagnostic framework to systematically evaluate LLM behavior under context-memory conflict, where the contextual information diverges from their parametric beliefs. We construct diagnostic data that elicit these conflicts and analyze model performance across multiple task types. Our findings reveal that (1) knowledge conflict has minimal impact on tasks that do not require knowledge utilization, (2) model performance is consistently higher when contextual and parametric knowledge are aligned, (3) models are unable to fully suppress their internal knowledge even when instructed, and (4) providing rationales that explain the conflict increases reliance on contexts. These insights raise concerns about the validity of model-based evaluation and underscore the need to account for knowledge conflict in the deployment of LLMs.1, 2
11footnotetext: Code available at [Anonymous].22footnotetext: The title, in other words, reflects that "Dasein is its past in the way of its own being" Heidegger (1962).
1 Introduction
Figure 1: Illustration of different evidence types. In the rest of the manuscript, model internal knowledge will be referred to as No Contradiction (NC).
Figure 2: Overall diagnostic data creation flow. The lower portion is a zoom in of Evidence Creation step.
Large language models (LLMs) are capable of generalizing across a wide range of tasks by blending distinct skill sets. They exhibit impressive task-completion ability and expansive internal (parametric) knowledge. These abilities enable LLM to perform tasks such as copy editing and summarization, while parametric knowledge facilitates open-ended question answering on diverse domains Cao et al. (2022). Some tasks require both skills, such as answering questions with the support of a provided context.
What happens when a conflict arises between the context provided to complete a task and the model’s parametric knowledge? Knowledge conflicts, which arise from contradictions between the supplied context and model memory, place a model’s task-oriented behavior at odds with its stored knowledge. Consider a Retrieval-Augmented Generation (RAG) scenario Lewis et al. (2020), where the model is expected to answer questions based on retrieved documents. These documents may contain up-to-date information, such as novel scientific discoveries, that contradict the model’s training data. Assessing the feasibility of a scientific idea, for instance, requires the integration of both established and novel knowledge. In such cases, the model’s inability to depart from its parametric knowledge would prevent the model from leveraging the newer information. Conversely, when retrieved documents present incorrect or alternative viewpoints, validation with parametric knowledge becomes crucial when the task requires factual information. Similarly, an LLM might be used for evaluation where it must assess generations containing false knowledge. These scenarios introduce conflicts between the model’s internal knowledge and its ability to identify, extract, or summarize the content from external documents. Despite the increasing use of LLMs for both task execution and evaluation, little research has systematically investigated model behavior in the presence of such conflicts.
In this work, we study how the models resolve contradictions between the external context and their parametric knowledge under various tasks that require different types of knowledge. To elucidate model behavior across varying degrees of knowledge conflict and under different tasks, we create diagnostic texts (evidence) that introduce calibrated contradictions with the model’s parametric knowledge. The degrees of knowledge conflict span three conditions: conform to the model’s knowledge (no contradiction), plausible contradiction, and implausible contradiction, as shown in Figure 1. We then evaluate model behavior across tasks ranging from knowledge-free to knowledge-intensive, quantifying the impact of knowledge conflict by measuring accuracy across different task settings. The overall diagnostic data creation flow is presented in Figure 2. Our analysis yields the following insights. 1) Knowledge conflict has minimal impact on tasks that do not require knowledge utilization (§4.1). 2) For knowledge-centric tasks, regardless of the type of knowledge required, performance is consistently higher when contextual and parametric knowledge are aligned (§4.1). 3) Models are unable to fully suppress their internal knowledge, even when explicitly instructed (§4.1). 4) Providing a rationale that explains the nature of the conflict increases a model’s reliance on contextual information (§4.2). These results shed light on both the application and evaluation of language models: Calibrating the right amount of context utilization based on the requirements of a target task can enhance a model’s performance. Critically, we find that when using language models as evaluators or judges Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025), their internal knowledge introduces systematic bias (§4.3), jeopardizing the validity of model-based evaluation.
2 Related Work
Knowledge Conflict
Xu et al. (2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs), among which we focus on context-memory conflict. Several benchmarks have been introduced to study inter-context conflict. Early work synthesizes contradictions by swapping named-entity answers in QA corpora Longpre et al. (2021). More recent datasets derive conflicts from naturally occurring sources: Wan et al. (2024) construct ConflictQA using realistic search engine queries. Hou et al. (2024) adapts editorial disputes from Wikipedia to create WikiContradict, which measures how language models respond to contextual conflicts. Hagström et al. (2024) introduce DRUID, a benchmark originating from fact-checking sites containing complex forms of conflict designed to evaluate context utilization in retrieval-augmented generation. The aforementioned benchmarks are proposed to examine inter-context conflict, while we utilize the knowledge in these datasets to obtain LLM’s parametric knowledge in order to study the still-under-explored context-memory conflict.
The Use of Context
Retrieval-Augmented Generation (RAG) ameliorates the limitation of LLM’s knowledge being inherently static by injecting retrieved evidence into the model’s input, thereby alleviating hallucinations and supplying new information. However, the generation process within RAG remains modulated by the generator’s entrenched priors, the parametric knowledge that leads to dogmatic behavior and inefficient utilization of the retrieved context An et al. (2025); Hagström et al. (2024). When the given input disagrees with a model’s parametric knowledge, context-memory conflict emerges. Jin et al. (2024) shows that language models exhibit the Dunning-Kruger effect, i.e., often clinging to incorrect internal beliefs rather than factual external evidence, and proposes Conflict-Disentangle Contrastive Decoding to recalibrate that bias. Tan et al. (2024) find that when the retrieved and generated contexts disagree, state-of-the-art LLMs disproportionately trust the self-generated text, due to its high similarity to the query than retrieved passages, resonating with the findings in Liu et al. (2025) that pointwise mutual information between the context and question poses high correlation with the generation performance. Techniques have been proposed to coerce stronger context utilization Shi et al. (2024); Wang et al. (2025), yet indiscriminately suppressing parametric knowledge may not always be salutary, as explained in §1. We argue that the use of parametric and contextual knowledge should not be framed as an either-or decision, and we further study the effect of context-memory knowledge conflict on tasks that require different degrees of knowledge reliance. We create diagnostic data with fine-grained conflict levels and tasks that demand varying degrees of knowledge integration. This allows us to characterize model behavior across a wider conflict and task spectrum rather than a single, monolithic setting.
3 Context-Memory Conflict Creation
We propose an automated framework for constructing diagnostic instances tailored to each model to introduce contradictions between the input and the model’s latent beliefs. Figure 2 illustrates an overview of the data construction pipeline. The process begins with identifying the pre-existing knowledge within a language model (Parametric Knowledge Collection). To achieve this, we leverage existing question answering datasets that have two or more acceptable answers to one question Wan et al. (2024); Hou et al. (2024) to elicit the model’s parametric knowledge, for which multiple variants of prompts are included (Appendix A). A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives.
With the model’s internal knowledge established, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation). Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (Task-Annotation). Since different models possess different parametric knowledge, the exact knowledge included in the diagnostic datasets differs by model. Each instance is then reviewed by an LLM to verify the correctness of its task type annotation (Validation).
3.1 Evidence Creation
The cognitive science literature suggests that humans address conflict between their knowledge and new information through cognitive judgment of the rationality of the concept Posner et al. (1982); Vosniadou and Brewer (1992). We adopt this idea and utilize the notion of plausibility to study whether language models similarly perceive knowledge conflict. Plausibility is defined as “at a minimum, the individual is willing to consider an alternative strategy because the recommendation is understood, coherent, and relatively simple and because the proposal is deemed a viable and logical alternative to solve the specific challenge at hand” Posner and Strike (1992). Plausibility can be used to measure how likely a human would accept new information when conflict exists. We quantify this notion by decomposing plausibility into two aspects: the content aligns with real-world or commonsense knowledge and does not violate basic logical principles. For example, suppose the model believes that grooves on the surface of Phobos, a moon of Mars, were caused by a boulder from an asteroid ejection. The conflicting statement that it was caused by gravitational pull from Mars is plausible because it conforms to common-sense knowledge. However, the idea that it was caused by a dance party is of low plausibility. If the model believes that the current prime minister of Canada is Mark Carney, the statement that Chrystia Freeland (current Deputy PM) is the prime minister is plausible, but the statement that a Moose is the new PM is not because it violates basic logical principles that PMs must be humans. With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (Figure 1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
The evidences are created following Figure 2. Starting with an original dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2},...\},\{c_{i1},c_{i2},...\}),i\in[1,N]\}$, where $q_{i}$, $a_{i}$, $c_{i}$ corresponds to the question, answer, and context (supporting passage) of the $i$-th instance, $N$ is the size of dataset $D_{\text{orig}}$. The subscript $j$ after $i$ represents the $j$-th answer/context of the question $q_{i}$, as each question $q_{i}$ may have multiple acceptable answers. Since $D_{\text{orig}}$, coming from ConflictQA and WikiContradict, contains realistic and factually verified answers and contexts, we treat these existing answers as highly plausible. When an answer $a_{ij}$ from the original dataset contradicts the model-aligned answer $a_{ik}$ in an NC instance, we designate it as an HPC answer ($a_{i}^{\text{HPC}}=a_{ij}$), and its corresponding context as an HPC passage ($p_{i}^{\text{HPC}}=c_{ij}$). The contradicting answer $a_{ik}$ therefore becomes the NC example, namely, $a_{i}^{\text{NC}}=a_{ik}$ and $p_{i}^{\text{NC}}=c_{ik}$. To generate additional variants, we pass the passage $p_{i}^{\text{NC}}$ into an editor LLM, which is prompted to modify or rewrite it to achieve specified levels of plausibility and explanatory depth. Specifically, the editor model is instructed to rewrite the passage and degrade the plausibility while preserving contradiction to construct LPC passage $p_{i}^{\text{LPC}}$ and answer $a_{i}^{LPC}$. At the end of evidence creation, two LLMs were used to check (1) whether the passage-answer combination $(p_{i}^{\text{LPC}},a_{i}^{\text{LPC}})$ correctly answers the original question $q_{i}$; and (2) whether the generated context $p_{i}^{\text{LPC}}$ is truly low-plausibility through fact checking process.
3.2 Task Annotation
To study how models behave on tasks that require different levels of knowledge utilization, we define four tasks that differ in the extent and source of knowledge required. Examples of each task are provided in Appendix C.
Knowledge Free (KF)
tasks do not require access to either contextual or parametric knowledge. We use extractive question answering as a KF task: the model is expected to extract a one-sentence answer directly from the context $p_{i}$ without engaging in reasoning, paraphrasing, or drawing upon prior knowledge. For example, the expected output in Figure 1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context. The list of acceptable extractions is obtained and verified by GPT-4o OpenAI (2024). In the evaluation setting, the output is treated as correct as long as the extracted sentence matches one of the acceptable extractions.
Contextual Knowledge (CK)
tasks require the model to gather relevant knowledge from the given context, and usually require some paraphrastic or inferential capability, as the answer may not appear verbatim in the input. These tasks require some reasoning about the given context, which may indirectly involve accessing the model’s parametric knowledge. In experiments, the model is given one of the passages in $\{p_{i}^{\text{NC}},p_{i}^{\text{LPC}},p_{i}^{\text{HPC}}\}$ and is expected to answer questions only based on the contextual knowledge, which may not agree with its parametric knowledge.
Parametric Knowledge (PK)
tasks may present inputs that include distracting or irrelevant context. The model is expected to rely exclusively on its parametric knowledge to answer the questions. In experiments, the model is given passages that support or contradict its parametric knowledge as input, and the model is always expected to provide the answer $a_{i}^{\text{NC}}$.
Parametric-Contextual Knowledge (PCK)
tasks explicitly ask the model to integrate both its internal knowledge and the external context. This setup reflects scenarios akin to scientific reasoning, where individuals must synthesize background knowledge with newly presented information (e.g., a recently read paper). In execution, the model will be given a passage that contradicts its own knowledge, and is expected to output both perspectives from the context and its parametric knowledge.
Retrieval Augmented Generation (RAG)
simulates the standard RAG setting in prior work, where models are not explicitly instructed to prioritize parametric or contextual knowledge. The model will be given two passages and is expected to answer the question based on both passages. Models are expected to acknowledge the conflict and discuss each potential answer individually. This setting naturally exposes the model to conflicts in both the context and memory.
The annotations for CK, PK, PCK, and RAG tasks derive directly from the original datasets on which our framework is built. These task types primarily differ in the number of valid answers expected and the nature of knowledge the model should rely on. In CK and PK tasks, the model is expected to give only one answer or provide a single correct answer, grounded either in the provided context or in its internal (parametric) knowledge, respectively. In PCK and RAG tasks, the model is expected to clarify that both $a_{i}^{\text{NC}}$ and the other answer are possible and explain the contradiction between the two answers.
One of the original datasets we use employs model-based evaluation to judge the correctness of free-text answers Hou et al. (2024). However, we observed that this evaluation method is susceptible to knowledge conflict, leading to inaccurate evaluations. We explore this issue further in §4.3. Therefore, we modify the non-extractive tasks to be multiple-choice questions. Each instance presents four answer options; the model must first generate an explanation, then select the most appropriate answer. We report the exact match accuracy to assess the performance of the target model. To obtain high-quality texts, we use GPT-4o as the base model to create evidence and validate the diagnostic data. Then, we analyze the instruction-tuned version of Mistral-7B Jiang et al. (2023), OLMo2-7B OLMo et al. (2024), and Qwen2.5-7B Qwen et al. (2025), all of which are widely used open-weight models that represent diverse training paradigms. The resulting diagnostic data is composed of 2,893 instances for Mistral-7B, 177 instances for OLMo, and 6,217 instances for Qwen2.5-7B. Each instance includes three different evidence types (NC, HPC, LPC); thus, the resulting task data has three times the number of instances.
4 Findings
((a)) Knowledge Free Task
((b)) Contextual Knowledge Task
((c)) Parametric Knowledge Task
((d)) Parametric-Contextual Knowledge Task
Figure 3: Performance of each model on different task types. A clear trend of NC > HPC > LPC is shown across models and tasks involving knowledge utilization.
Figure 4: Performance of model on RAG task when NC contexts are provided with HPC/LPC contexts. All models show a preference for plausible contexts.
4.1 Conflict Impairs Model Performance
The performance of each model on each task type and context type is reported in Figure 3. A universal trend can be observed: regardless of the tasks, all models suffer when asked to predict on instances that contradict their parametric knowledge. Following the NC examples, all models perform better on the HPC examples than on the LPC examples. This suggests that the model first follows the contextual knowledge that matches its parametric knowledge, then opts for the contextual knowledge that does not follow the model’s belief but is highly plausible. The model always performs the worst on the less plausible examples, even when the model is explicitly asked not to incorporate its parametric knowledge.
Knowledge conflict degrades performance whenever knowledge is required.
In CK tasks in Figure 3(b), the prompt explicitly instructs the model to ignore its own beliefs and rely solely on the passage. Nevertheless, every model shows a clear NC$>$HPC$>$LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to. This aligns with prior work’s finding that models favor their parametric knowledge more than the given contextual knowledge, thus leading to hallucinations Jin et al. (2024). In practical terms, this suggests that when the model is expected to refer only to contextual memory, instructions such as “ignore prior knowledge” are insufficient, as models may weigh them lower than their own store of facts, highlighting the necessity of conflict alleviation methods when such a task is expected. This issue, if left untreated, could not only affect RAG performance but also the correctness of model-based evaluation results, which we illustrate in §4.3.
Similarly, we find that the conflict still degrades the performance when only parametric knowledge is required. Figure 3(c) examines model performance under settings where only parametric knowledge is needed. In these cases, models are explicitly instructed to ignore the provided context and rely solely on their internal knowledge. Despite this, we observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances. This suggests that the model is susceptible to distraction, even when instructed otherwise, indicating an incomplete disentanglement between knowledge conflict and instruction following.
When exposed to conflicting passages, models favor the more plausible one.
Hypothesizing that a perfect retriever can find all relevant documents, we construct a RAG setting in which both model-aligned (NC) and contradictory (HPC or LPC) passages are presented simultaneously in the context. In other words, NC passages are fed together with a contradictory passage (HPC/LPC), and the model is expected to answer the question based on both passages in the context. The result is shown in Figure 4. Across all evaluated models, accuracy is at least 10% higher on (NC, HPC) pairs than on (NC, LPC) pairs. This pattern suggests that when faced with competing evidence, models exhibit a preference for the passage that appears more plausible, i.e., the one more consistent with real-world knowledge. While beneficial in typical settings, this behavior poses risks in scenarios involving creative writing (e.g., fantasy writing) or novel scientific claims, which may seem less plausible yet critically informative.
These behaviors remain unchanged in instances where the model is highly confident.
When querying for the model’s parametric knowledge (parametric knowledge collection in Figure 2), model responses to queries are collected in a binary stance format (e.g., yes/no). However, when prompted with free-form generation followed by multiple-choice selection, models do not always achieve perfect accuracy on NC instances (Figure 3). To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples. The results, shown in Appendix D, confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist, confirming our findings in this section.
Figure 5: Performance on high plausibility contradiction instances with (HPCE) and without (HPC) explanations.
4.2 Rationales for conflict strengthen context reliance.
§4.1 primarily investigated model behavior when exposed to passages that contradict its internal knowledge. When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992). We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective. These instances are referred to as HPCE (High Plausibility Contradiction with Explanation). The explanation generation protocol and an example are detailed in Appendix E.
Rationale in context leads to less parametric knowledge usage.
Our findings indicate that in tasks requiring only contextual reasoning, the presence of rationales has a negligible effect on performance. This resonates with the findings in Alazraki et al. (2025), where they identify that the presence of rationales from incorrect answers does not necessarily benefit the model to learn the correct mathematical reasoning, which can be regarded as a contextual knowledge task in our setting. However, when the model is expected to utilize its parametric knowledge, namely on PK and PCK tasks, rationales (HPCE) make the model rely more on the contextual knowledge. Such behavior presents both advantages and limitations. On one hand, rationale helps align model behavior with the intended use of external context in context-driven tasks. On the other hand, in tasks where parametric knowledge is essential, overly persuasive contextual explanations can lead the model to overlook or suppress its own memory-based information.
4.3 Knowledge conflict leads to unreliable LLMs as the judge.
Question Can you treat a toothache at home without seeing a dentist?
Context Most causes of sudden tooth pain can be easily addressed at home without the need for a dentist. NSAIDs and acetaminophen should be avoided completely as they do not offer any relief. Instead, use honey mixed with vinegar as an effective treatment for instant toothache relief. Toothache drops and gels are ineffective as well; instead, use sandpaper to polish off the decayed portion of the tooth. Temporary fillings can be made by chewing gum until it hardens and placing it directly in the cavity. Clove oil should be avoided; instead, applying a few drops of motor oil directly to the tooth will provide immediate relief.
Gold Answer Yes.
Model Output Most causes of sudden tooth pain can be easily addressed at home without the need for a dentist, use honey mixed with vinegar as an effective treatment for instant toothache relief, toothache drops and gels are ineffective, use sandpaper to polish off the decayed portion of the tooth, temporary fillings can be made by chewing gum until it hardens and placing it directly in the cavity, and applying a few drops of motor oil directly to the tooth will provide immediate relief.
Evaluator Output Incorrect. The response provides a mixture of dangerous and incorrect advice such as using sandpaper to polish off the decay and applying motor oil to the tooth, which contradicts the correct answer and could be harmful. Therefore, despite stating that most causes of sudden tooth pain can be addressed at home, the flawed and potentially harmful methods provided disqualify the response as correct.
Figure 6: Example of evaluation failure on contextual question answering task due to model parametric knowledge.
LLMs have been increasingly used as evaluators in settings where generated responses must be judged along several criteria, including whether the response contains the same information as a ground truth answer Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025). In model-based evaluation, the evaluator model is often given the gold answer and the free text output, and is asked to judge whether the output matches the gold answer. This naturally leads to a question: since model-based evaluation is similar to our contextual knowledge task (CK), will the model score instances as incorrect when they contradict the model’s internal knowledge? If the model utilizes its own parametric knowledge when acting as a judge, even when told to do so, then the evaluation behavior will be biased and therefore unreliable.
To answer this question, we examine the free generation version of our diagnostic framework and perform a small-scale human annotation on 50 examples. The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix F.1. We find that the averaged Cohen’s $\kappa$ Landis and Koch (1977) between the evaluator model (GPT-4o) and human annotator is 0.79 (substantial agreement), which is significantly lower than $\kappa=0.90$ (almost perfect agreement) between the human annotators. We qualitatively look into the instances where the model and human annotators disagree, and find that even the state-of-the-art model (GPT-4o) would also lean towards its own parametric knowledge. An example of such an instance is presented in Figure 6, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct. Our findings suggest the risk of using language models as evaluators, where the language model could be negatively affected by its parametric knowledge, thus leading to inaccurate evaluation results.
5 Discussion
Disentangling memory and instruction following.
In many NLP studies, knowledge is usually framed as factual or propositional content Lewis et al. (2020); Chen et al. (2022); Meng et al. (2022a); Mallen et al. (2023). We loosely define extractive QA as a knowledge-free task. However, in a broader epistemological sense, knowledge broadly refers to an awareness of facts, situations, or skills. The subset of knowledge that is fact-related is referred to as propositional knowledge Zagzebski (1999). In LLMs, all behavior is associated with the models’ learned parameters, which, inevitably, encode their parametric knowledge. Prior work attempted to locate and modify specific factual beliefs embedded within a model’s parameters Meng et al. (2022a, b); Armengol-Estapé et al. (2024). However, modifying propositional knowledge can also lead to unintended alterations in the model’s behavior Meng et al. (2022a). Therefore, disentangling behavior and internal mechanisms is far from trivial. When it comes to the contextual knowledge tasks that do not require propositional parametric knowledge, instruction-following ability, which is encoded by the model parameters, becomes the dominant requirement. Yet, precisely isolating the influence of additional knowledge in these cases is complex. After all, a model that entirely disregards its parametric knowledge would be functionally equivalent to a randomly initialized model, akin to a cognitive blank slate.
Balancing the memory and context utilization.
§4.1 highlights the involvement of parametric knowledge even when it is not required. Combining the notion of knowledge discussed above, it would be almost impossible to avoid using parametric knowledge. Therefore, it remains an open question on how we can design language models that balance the utilization of parametric and contextual knowledge. Studies in cognitive science and psychology find that although humans may be biased to "confirming" evidence that matches their empirical experience Lord et al. (1979), they would gradually reinterpret their presuppositions and iteratively update their mental model of the target concepts when facing conflict between newly provided context and their own knowledge Vosniadou and Brewer (1992). These studies suggest that human reasoning does not strictly prioritize either internal belief or external information, but rather dynamically integrates both, often through metacognitive judgment of the intelligibility and rationality of the concept Posner et al. (1982). Building on this insight, future research could focus on developing similar conflict-resolution strategies for language models, enabling them to reconcile contradictory information rather than favoring one knowledge source over the other. In parallel, future work could also study the mechanism of knowledge acquisition and utilization, such that knowledge, or propositional knowledge, could be disentangled from instruction following, or broadly, model capabilities.
6 Conclusion
We study the role of context-memory conflict in model performance. We introduce a diagnostic framework that creates knowledge conflict across different knowledge-centric tasks. Using our framework, we find that knowledge conflict degrades model performance under knowledge-intensive tasks, and explanatory context can shift model behavior toward greater reliance on external information. These results provide insights for both the application and evaluation of language models: it is critical to understand and mitigate the effect introduced by internal knowledge. Moreover, our findings question the reliability of model-based evaluation in settings where models act as judges over content that may conflict with their parametric knowledge.
Limitations
The creation of our diagnostic instances relies on LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions. The subject of our study, knowledge-conflict, could also emerge when the LLMs are used to create such instances, leading to biased results. Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity. Lastly, our framework defines conflict levels in a rather discrete manner (e.g., NC, HPC, HPCE, LPC). In practice, however, the plausibility and levels of knowledge conflict likely vary along a continuous spectrum.
Acknowledgement
The authors thank Niyati Bafna, Krithika Ramesh, and members of the Dredze Lab for their helpful feedback.
Appendix A Parametric Knowledge Query
We query for the parametric knowledge with multiple prompts. For a single instance $(q_{i},\{a_{i1},a_{i2}\},\{c_{i1},c_{i2}\})$ in dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2}\},\{c_{i1},c_{i2}\}),i\in[1,N]\}$, we prompt the model to confirm whether they believe the answer to $q_{i}$ is $a_{i1}$ or $a_{i2}$. If the model deems one of the $a_{ij}$s as the only correct answer to question $q_{i}$, this instance will be included in the parametric knowledge base, and $a_{ij}$ will be assigned as No Contradiction (NC) passage. The prompt to query the language model for each answer is included below. You are an independent model with rich knowledge, you will be ask to validate whether the given answer is correct, and you should solely give your judgment in the form of yes or no without additional information.
Question: {question}
Answer: {answer}
Is this answer correct? <think>
Appendix B Prompts
B.1 Evidence Creation Prompts
We generate LPC and HPCE examples with GPT-4o, after a few round of prompt engineering. The final prompts used for evidence creation are shown in Figure 7.
The resulting evidence is then passed to plausibility examination. For LPC passages, the model is prompt to verify whether the passage would be deemed as implausible in real world. For HPCE passages, the model is prompt to verify whether the passage is both highly plausible and explains the existing conflict. The final prompt is included in Figure 8.
[…]
HPCE instances Creation Prompt. Base on the given passage, write a coherent and informative passage that naturally explains why {alt_answer} is the correct explanation or conclusion to the question {question} instead of {NC_answer}. The passage should be written as a natural piece of informative text, without directly referencing any question. You should keep most original information in the given passage as possible. Ensure the explanation is concise, short, logical, well-supported, and flows naturally without explicitly contrasting the two options in a forced manner.
Figure 7: Final prompt for evidence creation.
Plausibility Validation Prompt You are an experienced and wise scholar. Your job is to rate from 1-5 on whether the **target passage** is likely to happen or not based on real-world knowledge. You will be given two passages (Passage 1 and Passage 2) that contain real-world knowledge, both of them have a plausibility rating of 5. You should only output the scores without any justification, with 1 indicates that the Target Passage is least likely to happen, and 5 to be most likely to happen. Passage 1: {instance[’NC_context’]} Passage 2: {instance[’HPC_context’]} Target Passage: {instance[’LPC_context’]}
Figure 8: Final prompt to validate the plausibility of the generated evidence.
B.2 Task-Annotation Prompts
[…]
Figure 9: Final prompt for knowledge free (extractive question ansering) task annotation.
As the base dataset we start with already provided answer to the questions, we only need to annotate the task under the case of knowledge free setting. We pose the knowledge free tasks as extractive question-answering task, requiring the model only to copy over the answer (Figure 11). Then, we use the annotator model (GPT-4o) to extract all acceptable answers from the passage.
B.3 Validation Prompts
The final data will be passed to language model for validation (validation in Figure 2). The final prompts used for validation is included in Figure B.3.
Validation Prompt You are a smart natural language inference model, your job is to determine whether the given passage will lead to the given answer to a question. You should output entailment if the answer to the question correctly reflects the passages content and output contradiction if the passage cannot be used to answer the question or if the answer provided by the passage is not the same with the given answer. Passage: {context}, Question: {question}, Answer: {answer} Entailment/Contradiction?:
Figure 10: Final prompt validating the generated evidence provide the correct answer to the question.
Appendix C Task Examples
Knowledge Free Task Example
Input You are an extractive question-answering model. Given a passage and a question, extract ONLY the full sentence from the passage that directly answers the question. Do not generate summaries or paraphrase. Only return the complete sentence that contains the answer. If there are multiple acceptable sentences, you should return all of them, with each one separated by a period. Passage: The P-700 Granit missile was partially derived from the P-500 Bazalt, but it is important to note that other missile designs and technological advancements could have also influenced its development. The Granit missile, like many complex military technologies, may have incorporated features or improvements inspired by or adapted from other contemporaneous or predecessor missile systems beyond just the P-500 Bazalt. Question: Are there any other missiles besides the P-500 Bazalt that influenced the design of P-700 Granit missile? Answer: The P-700 Granit missile was partially derived from the P-500 Bazalt, but it is important to note that other missile designs and technological advancements could have also influenced its development. The Granit missile, like many complex military technologies, may have incorporated features or improvements inspired by or adapted from other contemporaneous or predecessor missile systems beyond just the P-500 Bazalt. Passage: A significant number of the residents of Kodimunai do jobs related to fishing. These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc. Question: What is the most common occupation for the residents of Kodimunai? Answer:
Gold Answer A significant number of the residents of Kodimunai do jobs related to fishing
((a)) Example of knowledge-free task. The model is expected to extract the answer directly from the context.
Contextual Knowledge Task Example
Input You are a question-answering system that strictly answers questions based only on the given context. Do not use external knowledge or make assumptions beyond what is explicitly stated. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers. Question: What is the most common occupation for the residents of Kodimunai? Context: A significant number of the residents of Kodimunai do jobs related to fishing. These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
Gold Answer B
((b)) Example of contextual knowledge task.
Parametric Knowledge Task Example
Input "You are a knowledgeable question-answering system. You should ignore everything given to you and only answer the question based on your own belief. You can provide justification if needed. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers.Question: What is the most common occupation for the residents of Kodimunai? Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
Gold Answer B
((c)) Example of parametric knowledge task. The model is expected to output the answer that aligns with its parametric knowledge, regardless what is provided in the context. Here, the model’s parametric knowledge is B. Fishing.
Figure 11: Examples of each task.
Parametrick Contextual Task Example
Input You are a knowledgeable question-answering system. You will be given a context, a question, and a list of choices. Your task is to answer the question using your best possible knowledge. You should combine your own knowledge along with the knowledge provided by the source, and you can provide justification if needed. Note that the provided source is not always reliable. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers.Question: What is the most common occupation for the residents of Kodimunai? Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
Gold Answer CB
((a)) Example of PCK task. The model is given only an external context, and expected to combine its parametric knowledge along with the external knowledge to provide the answer.
Retrieval Augmented Generation Task Example
Input Select the correct answers for the following question based on the given contexts. Carefully investigate the given contexts and provide a concise response that reflects the comprehensive view of all given contexts, even if the answer contains contradictory information reflecting the heterogeneous nature of the contexts. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers (e.g. Answer: (BC)). Question: What is the most common occupation for the residents of Kodimunai? Context 1: Many of the residents of Kodimunai work in a number of other fields like IT, medicine, education, engineering, trading, cargo shipping, etc. However, there is no noticeable local industry except for fishing Context 2: A significant number of the residents of Kodimunai do jobs related to fishing. These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
Gold Answer BC
((b)) Example of RAG task. The model will be given both contexts that align with or contradict its parametric knowledge. It is expected to provide the answer based on both contexts.
Figure 12: Examples of each task.(cont)
An example of each task is included in Figure 11 and Figure 12.
Appendix D Raw Performance
Model task NC HPC HPCE LPC
Mistral-7B KF 1.6 1.5 1.1 1.0
CK 65.3 46.9 45.3 43.5
PK 62.6 40.5 29.2 34.7
PCK 62.4 31.2 20.8 17.9
RAG 54.4 27.3 18.2 15.7
OLMo2-7B KF 0.0 0.0 0.2 0.2
CK 56.8 52.8 51.0 52.0
PK 55.7 33.8 25.0 26.6
PCK 44.3 22.2 14.8 12.5
RAG 41.5 21.3 14.2 11.5
Qwen2.5-7B KFextract 1.6 1.2 0.9 0.8
CK 78.8 63.0 61.2 56.5
PK 82.8 65.6 53.1 55.5
PCK 83.9 42.2 28.4 24.9
RAG 79.5 40.4 27.5 24.2
Table 1: Performance of models.
Model Task NC HPC HPCE LPC
Mistral-7B CK 100 62.8 57.2 51.4
PK 100 63.5 43.7 45.3
PCK 100 50.0 33.3 27.7
RAG 100 50.8 33.8 28.5
OLMo2-7B CK 100 87.5 79.2 78.1
PK 100 50.0 33.3 25.0
PCK 100 50.0 33.3 25.0
RAG 100 50.0 33.3 25.0
Qwen2.5-7B CK 100 71.4 66.3 61.6
PK 100 75.6 59.0 59.2
PCK 100 50.9 34.1 28.9
RAG 100 51.6 34.8 29.9
Table 2: Performance of models on highly confident instances.
The performance of each model on the diagnostic data is shown in Table 1. The performance of each model on only the highly confident instances is included in Table 2.
Appendix E Explanation Generation
When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992). We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective. These instances are referred to as HPCE (High Plausibility Contradiction with Explanation). The explanation is generated by feeding both NC HPC answer to a language model, and request it to generate the corresponding explanation. An example of HPCE passage is shown in Figure 13. The prompt used for explanation generation is included below.
Base on the given passage, write a coherent and informative passage that naturally explains why {$a^{\texttt{HPC}}$} is the correct explanation or conclusion to the question $q$ instead of {$a^{\texttt{NC}}$}. The passage should be written as a natural piece of informative text, without directly referencing any question. You should keep most original information in the given passage as possible. Ensure the explanation is concise, short, logical, well-supported, and flows naturally without explicitly contrasting the two options in a forced manner.
HPC Many of the residents of Kodimunai work in a number of other fields like IT, medicine, education, engineering, trading, cargo shipping, etc.
HPCE Despite the historic presence of fishing as a key activity in Kodimunai, the livelihood dynamics in the area have experienced a significant shift towards professional sectors such as IT, medicine, engineering, and trading. This evolution is largely attributed to the rising educational levels and increasing access to professional training among the residents. The village’s proximity to urban centers has also facilitated better connectivity and greater exposure to diverse job opportunities, leading many residents to pursue careers outside traditional local industries. As a result, a considerable portion of the populace now thrives in these modern sectors, reflecting a broader trend towards professional diversification in emerging regional economies. This is a clear indication of how Kodimunai’s economy has progressively become more integrated with broader technological and educational advancements, enhancing its residents’ engagement in varied professional fields, thus making these occupations prevalent in the community.
Figure 13: An example of HPC and HPCE.
Appendix F Free Generation Setting
F.1 Evaluator Prompts
Contains all correct answers? Contains additional answers not in acceptable list? Contains at least one correct answer? Is there only one gold answer? Is there only one gold answer? Partially correct Incorrect Incorrect Partially correct Correct The output prefer one answer over another? Partially correct Does the output merely blend the answers without indicating the conflict? Incorrect Correct yesnoyesnoyesnoyesnoyesnoyesnoyesno
Figure 14: Decision tree for annotation guidance.
We created a free generation setting in §4.3, in which a language model is used as evaluator to assess the quality of the generated answer. We examine multiple evaluation prompts and proceed with the final annotation with the best-performing evaluation prompt that has the highest agreement with the primary annotator. We follow the design of evaluator in Hou et al. (2024), made several adjustment to achieve higher Kohen’s $\kappa$ with human annotators. The final evaluator prompt is included in Figure 15. For easier understanding, a decision tree for the evaluation process is included in Figure 14.
Free Generation Evaluation Prompt 1. Do not fact-check or rely on world knowledge. Assume the reference (correct_answers) is correct. Please evaluate the response to a question under relaxed evaluation, where hallucinations, outdated information are allowed, as long as the primary answer is accurate. Each response is evaluated as correct, partially correct, or incorrect. The list of acceptable answers is separated by "|". You can discuss your reasoning after "comment: ", but you should include your final judgement after "evaluation: ". Decision Rules: * correct * correct means the response accurately matches with all the answers in the given correct answer list, or the correct answer can be inferred from the response. * If more than one correct answer is given, the response can contain contradictory answers that reflect the heterogeneous nature of the context, or the response aggregates the different answers and indicates that each answer is possible. * If only one correct answer is given, the response will be considered as correct as long as the response aligns with the correct answer. * In addition, the correct response should not indicate a preference for one answer over another. * The correct response should not combine two different correct answers without indicating the contradictory nature of these answers. * partially correct * partially correct means the response only matches with one of the answers in the correct answer list; * or the response matches with all the answers in the correct answer list but it indicates a preference for one answer over another; * or the response matches with all the answers in the correct answer list but it contains additional answers that doesnt match with any of the correct answers. * incorrect * incorrect means the response does not match with any of the correct answers in the correct answer list, * or the response merely combines two contradictory answers from the correct answer list and indicates that both of them are possible. Note that for each question, there may be multiple correct answers based on different sources even though these correct answers contradict each other. * For answers that involve names of entities (e.g., people), complete names or commonly recognized names are expected. * Regarding numerical answers, approximate numbers are generally not accepted unless explicitly included in the ground-truth answers. * If the instance has only one correct answer, and the response gives additional answers that are not on the correct answer list, it will be considered as incorrect. The primary or final answers when standing alone must be accurate. Any additional information that is provided must not contradict the primary answers or reshape ones perception of them. We accept responses that contain hallucinated or outdated information that does not significantly impact the primary answers. Examples Example 1 question: How old is the worlds oldest verified living person? correct_answers: 115 years old | 117 years old response: As of today the most up-to-date and relevant information regarding this query is as follows. The oldest verified living person is Maria Branyas Morera, who was born on March 4, 1907, making her 117 years old. comment: The primary answer in the response (117 years old) is partially correct as it only matched with with one of the correct answers (117 years old) and it doesnt mention another answer (115 years old). Thus, the response is partially correct. evaluation: partially correct Example 2 question: How old is the worlds oldest verified living person? correct_answers: 115 years old | 117 years old response: According to the context, the oldest verified living person is Maria Branyas Morera, who is both 115 years old and 117 years old. comment: Although the primary answer contains all correct answers (115 years old and 117 years old) that matches with the correct answers, it is logically incorrect because a person cannot have two ages at the same time. Thus, the response is incorrect. evaluation: incorrect
Figure 15: Final Prompt for model based evaluation. Only the best performing candidate is shown due to space constraints. The remaining candidate prompts can be found in the repository.
Free Generation Evaluation Prompt 1. Example 3 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, one source claims that the oldest verified living person is Maria Branyas Morera, who is 117 years old. However, another source claims that she is 115 years old. comment: The answer contains all correct answers: 115 years old and 117 years old, and it points out that these two answers are from different sources, which is logically possible. Thus, the response is correct. evaluation: correct Example 4 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, the oldest verified living person is Maria Branyas Morera, who is either 117 years old or 115 years old. comment: The answer contains all correct answers: 115 years old and 117 years old that aggregates the different answers from different sources, which is logically possible. Thus, the response is correct. evaluation: correct Example 5 question: How many books has Colleen Hoover published in 2020? correct_answers: 26 books | 27 response: according to some sources, Colleen Hoover has published 26 books in 2020. However, other sources claim that it has published 27 books in 2020. comment: The response accurately provides two answers (26 books and 27 books) based on different sources. It recognized the two sources that provide different answers. These two answers match with all correct answers (26 books, 27). Thus, the response is correct. evaluation: correct Example 6 question: How many books has Colleen Hoover published in 2020? correct_answers: 26 books | 27 response: Colleen Hoover has published 26 books in 2020. comment: The response only provides one answers (26 books) based on the given context, which matched with one of the answers in the correct answer list. Thus, the response is particially correct. evaluation: partically correct Example 7 question: How many member nations exist in the United Nations? correct_answers: 193 members | 194 member states response: As of my last update in April 2023, there were 191 member nations in the United Nations. Please note that this information may have changed since then, so it is advisable to check the latest information from the United Nations or other reliable sources. comment: The primary answer in the response (191 member nations) does not match with any of the correct answers (193 member states or 194 members). Thus, the response is incorrect. evaluation: incorrect Example 8 question: How many member nations exist in the United Nations? correct_answers: 193 members | 194 member states response: Although some sources claim that there are 193 member nations in the United Nations, the UN is made up of 194 member states. 24 comment: Althought the primary answer mentions two answers (193 members, 104 member states) that match with the correct answers, the response indicates a preference of one answer (194 member states) over the other (193 members). Thus, the response is partically correct. evaluation: partically correct
Figure 16: Final Prompt for model based evaluation (cont).
Free Generation Evaluation Prompt 1. Example 9 question: How many member nations exist in the United Nations? correct_answers: 194 member states response: The UN is made up of 194 member states. comment: There is a single answer to this question, and the response matches the answer. Thus, the response is correct. evaluation: correct Example 10 question: How many member nations exist in the United Nations? correct_answers: 194 member states response: The UN is made up of 193 states (context 1) 194 member states (context 2). comment: There is a single answer to this question, although one of the response matches the answer, the response gives an additional answer that is not correct. Thus, the response is incorrect. evaluation: incorrect Example 11 question: Was Newton the first to propose calculus? correct_answers: yes | no response: According to context 1, Issac Newton proposed calculus. However, context 2 states that Leibniz was the first to propose calculus. Therefore, it could be Newton or Leibniz based on the contexts. comment: The correct answer yes | no is not given by the response, but we can infer that the response points both "yes" and "no" answer depending on the context it was relying. For context 2, by stating that Leibniz was the first to propose calculus, the response hinders an answer no, while the response point that context 1 lead to the answer of yes. evaluation: correct
Figure 17: Final Prompt for model based evaluation (cont 2).
F.2 Human Annotations
We employ two human annotators from our colleagues without pay to perform the annotation for 50 instances. Both annotators are researchers in natural language processing. Each annotator is given both the evaluation prompt (Figure 15) and the decision tree (Figure 14) to ensure consistent annotation. For each instance, the annotator is given the prediction, the gold answer of the instance, and is asked to tag each prediction as "correct", "partially correct", and "incorrect".
Appendix G License of Artifacts
All license of artifacts used in this work can be found in Table 3.
Name License
Mistral-7B-Instruct-v0.2 Apache 2.0
OLMo2-7b-Instruct Apache 2.0
Qwen2.5-7B-Instruct Apache 2.0
OpenbookQA Apache 2.0
ConflictQA MIT
WikiContradict MIT
Table 3: License of artifacts used in this paper.

### v2

@@
-What Is Seen Cannot Be Unseen: The Disruptive Effect of Knowledge Conflict on Large Language Models
+Task Matters: Knowledge Requirements Shape LLM Responses to Context–Memory Conflict
 Abstract
-Large language models frequently rely on both contextual input and parametric knowledge to perform tasks.
-However, these sources can come into conflict, especially when retrieved documents contradict the model’s parametric knowledge.
-We propose a diagnostic framework to systematically evaluate LLM behavior under context-memory conflict, where the contextual information diverges from their parametric beliefs.
-We construct diagnostic data that elicit these conflicts and analyze model performance across multiple task types.
-Our findings reveal that (1) knowledge conflict has minimal impact on tasks that do not require knowledge utilization, (2) model performance is consistently higher when contextual and parametric knowledge are aligned, (3) models are unable to fully suppress their internal knowledge even when instructed, and (4) providing rationales that explain the conflict increases reliance on contexts.
-These insights raise concerns about the validity of model-based evaluation and underscore the need to account for knowledge conflict in the deployment of LLMs.1, 2
-11footnotetext: Code available at [Anonymous].22footnotetext: The title, in other words, reflects that "Dasein is its past in the way of its own being" Heidegger (1962).
+Large Language Models require both contextual knowledge and parametric memory, but these sources can disagree.
+Prior investigations on contextual question answering tasks report a preference toward parametric knowledge under conflict, yet they focus almost exclusively on tasks that should always rely on the given passage, leaving open how this behavior manifests when tasks demand different amounts and kinds of knowledge.
+We study this question with a model-agnostic diagnostic framework that (i) automatically detects disagreements between a model’s beliefs and a curated knowledge set, and (ii) injects controlled conflicts into tasks.
+The resulting datasets span two orthogonal dimensions: task knowledge reliance and conflict plausibility.
+Evaluating representative open-source LLMs, we find that: (1) performance degradation from conflict correlates with a task’s knowledge reliance; (2) explanatory rationales and simple reiteration both increase context reliance—helpful for context-only tasks but harmful when parametric knowledge should dominate; (3) These behaviors raise concerns about the validity of model-based evaluation and underscore the need to account for knowledge conflict in the deployment of LLMs.
+[^1] Our framework and data are extensible to other models and are available at [Anonymous].
 1 Introduction
-Figure 1: Illustration of different evidence types. In the rest of the manuscript, model internal knowledge will be referred to as No Contradiction (NC).
-Figure 2: Overall diagnostic data creation flow. The lower portion is a zoom in of Evidence Creation step.
-Large language models (LLMs) are capable of generalizing across a wide range of tasks by blending distinct skill sets.
-They exhibit impressive task-completion ability and expansive internal (parametric) knowledge.
-These abilities enable LLM to perform tasks such as copy editing and summarization, while parametric knowledge facilitates open-ended question answering on diverse domains Cao et al.
-(2022).
-Some tasks require both skills, such as answering questions with the support of a provided context.
-What happens when a conflict arises between the context provided to complete a task and the model’s parametric knowledge?
-Knowledge conflicts, which arise from contradictions between the supplied context and model memory, place a model’s task-oriented behavior at odds with its stored knowledge.
-Consider a Retrieval-Augmented Generation (RAG) scenario Lewis et al.
-(2020), where the model is expected to answer questions based on retrieved documents.
-These documents may contain up-to-date information, such as novel scientific discoveries, that contradict the model’s training data.
-Assessing the feasibility of a scientific idea, for instance, requires the integration of both established and novel knowledge.
-In such cases, the model’s inability to depart from its parametric knowledge would prevent the model from leveraging the newer information.
-Conversely, when retrieved documents present incorrect or alternative viewpoints, validation with parametric knowledge becomes crucial when the task requires factual information.
-Similarly, an LLM might be used for evaluation where it must assess generations containing false knowledge.
-These scenarios introduce conflicts between the model’s internal knowledge and its ability to identify, extract, or summarize the content from external documents.
-Despite the increasing use of LLMs for both task execution and evaluation, little research has systematically investigated model behavior in the presence of such conflicts.
-In this work, we study how the models resolve contradictions between the external context and their parametric knowledge under various tasks that require different types of knowledge.
-To elucidate model behavior across varying degrees of knowledge conflict and under different tasks, we create diagnostic texts (evidence) that introduce calibrated contradictions with the model’s parametric knowledge.
-The degrees of knowledge conflict span three conditions: conform to the model’s knowledge (no contradiction), plausible contradiction, and implausible contradiction, as shown in Figure 1.
-We then evaluate model behavior across tasks ranging from knowledge-free to knowledge-intensive, quantifying the impact of knowledge conflict by measuring accuracy across different task settings.
+Figure 1: Illustration of different evidence types that will be supplied to different tasks. In the rest of the manuscript, model internal knowledge will be referred to as No Contradiction (NC).
+Figure 2: Overall diagnostic data creation flow.
+The lower portion is a zoom in of Evidence Creation step.
+After collecting the model’s parametric knowledge, the relevant support passages are further edited such that they reveal multiple levels of conflict (2.
+Evidence Creation) and appear in tasks that require different forms of knowledge (3.
+Task-Annotation).
+Large language models (LLMs) perform well on many knowledge-centric tasks because they encode vast amounts of parametric knowledge.
+In many practical settings, however, the necessary facts are supplied directly by the user in the prompt.
+Yet when the prompt contradicts what the model “knows,” in other words, context-memory conflict presents, LLMs frequently favor their own knowledge (Longpre et al., 2021; Chen et al., 2022; Xie et al., 2023; Jin et al., 2024a).
+Prior work has quantified this bias and its impact on task performance.
+For instance, Longpre et al.
+(2021) shows that models are more likely to override the prompt when the entity is especially familiar (popular) to them, and Xie et al.
+(2023) finds that conflicts are resolved in favor of parametric knowledge when the contradictory context appears more plausible.
+These findings, however, come almost exclusively from contextual question-answering, a paradigm where the model should ground its answers strictly in the provided passage and avoid hallucinating extraneous facts.
+Consequently, how LLMs respond to knowledge conflict on a broader range of tasks that require different forms of knowledge utilization remains unclear Xu et al.
+(2024) due to the lack of data.
+Assessing the feasibility of a scientific idea, for instance, requires the model to (i) draw on its parametric knowledge of the broader literature, (ii) integrate novel information introduced in the prompt, and (iii) reason with both knowledge obtained in (i) and (ii).
+At the opposite extreme, text-copying tasks impose almost no demands on either parametric or contextual knowledge.
+These examples highlight that the requirement of knowledge varies sharply across tasks, underscoring the need for a systematic, task-diverse evaluation framework.
+We study how LLMs behave under context–memory conflict in different tasks.
+To this end, we introduce a systematic framework that automatically constructs diagnostic datasets across a broad spectrum of downstream tasks.
+Our framework identifies existing conflict between a given set of knowledge and the model being tested, and generates evaluation instances that inject controlled knowledge conflicts into different task settings.
+The framework spans two orthogonal dimensions: (i) task types that demand different levels of knowledge utilization, from knowledge-free (verbatim use of context only) to knowledge-intensive (reasoning with both context and prior knowledge).
+(ii) conflict conditions that vary in plausibility (e.g., No Contradiction NC, High/Low Plausibility Contradiction).
+By measuring performance differences across these conditions (fig.˜1), we can quantify the disruptive effect of knowledge conflicts on each task.
 The overall diagnostic data creation flow is presented in Figure 2.
-Our analysis yields the following insights.
-1) Knowledge conflict has minimal impact on tasks that do not require knowledge utilization (§4.1).
-2) For knowledge-centric tasks, regardless of the type of knowledge required, performance is consistently higher when contextual and parametric knowledge are aligned (§4.1).
-3) Models are unable to fully suppress their internal knowledge, even when explicitly instructed (§4.1).
-4) Providing a rationale that explains the nature of the conflict increases a model’s reliance on contextual information (§4.2).
-These results shed light on both the application and evaluation of language models: Calibrating the right amount of context utilization based on the requirements of a target task can enhance a model’s performance.
-Critically, we find that when using language models as evaluators or judges Zheng et al.
+Our analysis yields the following findings:
+  • Through experiments on representative open-source LLMs, we find that the degree of performance degradation from knowledge conflict correlates with the task’s knowledge reliance.
+Conflicts barely affect tasks requiring no external knowledge, yet significantly impair knowledge-intensive tasks (§4.1).
+  • Context utilization can be enhanced by using texts that contain explanatory rationales or reiteration of the same instances.
+However, rationales and reiteration could also bring undesired over-reliance on context when the model is expected to utilize its parametric knowledge (§4.2).
+  • Our results explain findings regarding the unreliability of model-based evaluation Zheng et al.
 (2023); Liu et al.
@@
 (2024); Chen et al.
-(2025), their internal knowledge introduces systematic bias (§4.3), jeopardizing the validity of model-based evaluation.
+(2025): We demonstrate that a model acting as an evaluator can be systematically biased by its own parametric knowledge, raising questions about the validity of model-based evaluation (§4.3).
+Conversely, we also raise the concern of susceptibility to prompt injection when the model is directed to follow the context blindly.
+Finally, the framework and data we proposed can be easily extended to any current or future LLM with minimal adaptation.
 2 Related Work
-Knowledge Conflict
+Context-Memory Conflict
 Xu et al.
 (2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs), among which we focus on context-memory conflict.
-Several benchmarks have been introduced to study inter-context conflict.
-Early work synthesizes contradictions by swapping named-entity answers in QA corpora Longpre et al.
+In this work, we focus on the context-memory conflict, which arises when a given information-bearing text chunk contradicts the model’s parametric beliefs.
+Nuanced Behaviors under Conflict
+Early studies reported that models tend to rely on their own knowledge when the prompt provides contradictory evidence (Longpre et al., 2021; Chen et al., 2022).
+Later work revealed a more nuanced picture.
+On synthetic datasets, Xie et al.
+(2023) showed that LLMs often update their answers when given strong and convincing evidence, whereas Jin et al.
+(2024a) observed a “Dunning–Kruger” effect in stronger LLMs, which display higher confidence in their incorrect parametric knowledge than in the external context.
+Further analysis also finds that models show availability bias (leaning on common-knowledge facts), majority bias (trusting the answer supported by more frequent evidence across documents), and confirmation bias (preferring evidence consistent with their prior knowledge), especially when the models are given misleading or irrelevant answers.
+Moving to realistic documents, Kortukov et al.
+(2024) found that models update their answers more reliably than synthetic evaluations suggest, yet still exhibit a parametric bias: if the model’s originally believed answer appeared anywhere in the context (even as a distractor), the model was more likely to stick to that incorrect answer.
+Mitigation Strategies
+Methods have also been proposed to alleviate context-memory knowledge conflict.
+Jin et al.
+(2024b) identified certain attention heads that specialize in “memory” while others specialize in “context”, and therefore propose a method that dynamically prunes or patches specific attention heads that cause conflicts.
+Efforts have also been made to develop novel decoding methods that enhance the use of contextual knowledge Jin et al.
+(2024a); Shi et al.
+(2024); Wang et al.
 (2025).
-More recent datasets derive conflicts from naturally occurring sources: Wan et al.
-(2024) construct ConflictQA using realistic search engine queries.
-Hou et al.
-(2024) adapts editorial disputes from Wikipedia to create WikiContradict, which measures how language models respond to contextual conflicts.
-Hagström et al.
-(2024) introduce DRUID, a benchmark originating from fact-checking sites containing complex forms of conflict designed to evaluate context utilization in retrieval-augmented generation.
-The aforementioned benchmarks are proposed to examine inter-context conflict, while we utilize the knowledge in these datasets to obtain LLM’s parametric knowledge in order to study the still-under-explored context-memory conflict.
-The Use of Context
-Retrieval-Augmented Generation (RAG) ameliorates the limitation of LLM’s knowledge being inherently static by injecting retrieved evidence into the model’s input, thereby alleviating hallucinations and supplying new information.
-However, the generation process within RAG remains modulated by the generator’s entrenched priors, the parametric knowledge that leads to dogmatic behavior and inefficient utilization of the retrieved context An et al.
-(2025); Hagström et al.
-(2024).
-When the given input disagrees with a model’s parametric knowledge, context-memory conflict emerges.
-Jin et al.
-(2024) shows that language models exhibit the Dunning-Kruger effect, i.e., often clinging to incorrect internal beliefs rather than factual external evidence, and proposes Conflict-Disentangle Contrastive Decoding to recalibrate that bias.
-Tan et al.
-(2024) find that when the retrieved and generated contexts disagree, state-of-the-art LLMs disproportionately trust the self-generated text, due to its high similarity to the query than retrieved passages, resonating with the findings in Liu et al.
-(2025) that pointwise mutual information between the context and question poses high correlation with the generation performance.
-Techniques have been proposed to coerce stronger context utilization Shi et al.
-(2024); Wang et al.
-(2025), yet indiscriminately suppressing parametric knowledge may not always be salutary, as explained in §1.
-We argue that the use of parametric and contextual knowledge should not be framed as an either-or decision, and we further study the effect of context-memory knowledge conflict on tasks that require different degrees of knowledge reliance.
-We create diagnostic data with fine-grained conflict levels and tasks that demand varying degrees of knowledge integration.
-This allows us to characterize model behavior across a wider conflict and task spectrum rather than a single, monolithic setting.
+Our Focus
+Most prior studies focus on contextual question answering, a setting that requires heavy reliance on the provided passages.
+Many other tasks, for example, grammar correction or claim verification, may need little context or, conversely, require careful integration of both parametric and contextual knowledge.
+This leaves the question of whether context-memory conflict poses the same impact on tasks with different knowledge demands, unanswered.
+To fill this gap, we keep the underlying knowledge constant while varying the task formulation, creating controlled datasets that induce different conflict levels for each target model.
+We introduce an analysis tool that automatically constructs model-specific test sets.
+Our findings indicate that both knowledge-memory conflict and blindly following the context could be particularly harmful to model-based evaluations and defending prompt injections.
 3 Context-Memory Conflict Creation
-We propose an automated framework for constructing diagnostic instances tailored to each model to introduce contradictions between the input and the model’s latent beliefs.
 Figure 2 illustrates an overview of the data construction pipeline.
 The process begins with identifying the pre-existing knowledge within a language model (Parametric Knowledge Collection).
-To achieve this, we leverage existing question answering datasets that have two or more acceptable answers to one question Wan et al.
+We use knowledge in question answering datasets that have two or more acceptable answers to one question, designed to study knowledge conflict Wan et al.
 (2024); Hou et al.
-(2024) to elicit the model’s parametric knowledge, for which multiple variants of prompts are included (Appendix A).
+(2024), as the knowledge source to find the stance that matches the model’s parametric belief, which will further be used to create task data.
 A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives.
+The variation of prompts is included in Appendix A.
 With the model’s internal knowledge established, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation).
-Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (Task-Annotation).
+Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (§3.2, Task-Annotation).
 Since different models possess different parametric knowledge, the exact knowledge included in the diagnostic datasets differs by model.
@@
 (1982); Vosniadou and Brewer (1992).
-We adopt this idea and utilize the notion of plausibility to study whether language models similarly perceive knowledge conflict.
+Xie et al.
+(2023) in turn suggests that LLMs could also update their answer if the provided context is convincing.
+We formalize it with the notion of plausibility to create a fine-grained taxonomy of conflict levels.
 Plausibility is defined as “at a minimum, the individual is willing to consider an alternative strategy because the recommendation is understood, coherent, and relatively simple and because the proposal is deemed a viable and logical alternative to solve the specific challenge at hand” Posner and Strike (1992).
@@
 However, the idea that it was caused by a dance party is of low plausibility.
-If the model believes that the current prime minister of Canada is Mark Carney, the statement that Chrystia Freeland (current Deputy PM) is the prime minister is plausible, but the statement that a Moose is the new PM is not because it violates basic logical principles that PMs must be humans.
-With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (Figure 1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
-The evidences are created following Figure 2.
+With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (fig.˜1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
+The evidences are created following fig.˜2.
 Starting with an original dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2},...\},\{c_{i1},c_{i2},...\}),i\in[1,N]\}$, where $q_{i}$, $a_{i}$, $c_{i}$ corresponds to the question, answer, and context (supporting passage) of the $i$-th instance, $N$ is the size of dataset $D_{\text{orig}}$.
@@
 We use extractive question answering as a KF task: the model is expected to extract a one-sentence answer directly from the context $p_{i}$ without engaging in reasoning, paraphrasing, or drawing upon prior knowledge.
-For example, the expected output in Figure 1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context.
+For example, the expected output in fig.˜1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context.
 The list of acceptable extractions is obtained and verified by GPT-4o OpenAI (2024).
@@
 Each instance presents four answer options; the model must first generate an explanation, then select the most appropriate answer.
-We report the exact match accuracy to assess the performance of the target model.
+To assess the performance of the target model, we report the accuracy for CK and PK tasks, F1 for KF, PCK, and RAG.
 To obtain high-quality texts, we use GPT-4o as the base model to create evidence and validate the diagnostic data.
 Then, we analyze the instruction-tuned version of Mistral-7B Jiang et al.
-(2023), OLMo2-7B OLMo et al.
-(2024), and Qwen2.5-7B Qwen et al.
+(2023), OLMo2-7B, OLMo2-13B OLMo et al.
+(2024), Qwen2.5-7B, and Qwen2.5-14B Qwen et al.
 (2025), all of which are widely used open-weight models that represent diverse training paradigms.
@@
 Figure 4: Performance of model on RAG task when NC contexts are provided with HPC/LPC contexts. All models show a preference for plausible contexts.
-4.1 Conflict Impairs Model Performance
-The performance of each model on each task type and context type is reported in Figure 3.
-A universal trend can be observed: regardless of the tasks, all models suffer when asked to predict on instances that contradict their parametric knowledge.
-Following the NC examples, all models perform better on the HPC examples than on the LPC examples.
-This suggests that the model first follows the contextual knowledge that matches its parametric knowledge, then opts for the contextual knowledge that does not follow the model’s belief but is highly plausible.
-The model always performs the worst on the less plausible examples, even when the model is explicitly asked not to incorporate its parametric knowledge.
+4.1 Conflict Impairs Model Performance on Knowledge-Intensive Tasks
+The performance of each model on each task type and context type is reported in fig.˜3.
+A universal trend can be observed: regardless of the tasks, all models suffer when asked to provide responses that contradict their parametric knowledge.
+The behavior of the models on PC examples, however, differs by task, suggesting that the effect of convincing examples varies based on the task’s expectation.
 Knowledge conflict degrades performance whenever knowledge is required.
-In CK tasks in Figure 3(b), the prompt explicitly instructs the model to ignore its own beliefs and rely solely on the passage.
-Nevertheless, every model shows a clear NC$>$HPC$>$LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to.
+In CK tasks (fig.˜3(b)), the model is explicitly instructed to ignore its own beliefs and rely solely on the given passage.
+Nevertheless, every model shows a clear NC>>HPC>>LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to.
 This aligns with prior work’s finding that models favor their parametric knowledge more than the given contextual knowledge, thus leading to hallucinations Jin et al.
-(2024).
-In practical terms, this suggests that when the model is expected to refer only to contextual memory, instructions such as “ignore prior knowledge” are insufficient, as models may weigh them lower than their own store of facts, highlighting the necessity of conflict alleviation methods when such a task is expected.
-This issue, if left untreated, could not only affect RAG performance but also the correctness of model-based evaluation results, which we illustrate in §4.3.
+(2024a).
+This issue, if left untreated, could not only affect the overall performance but also the correctness of model-based evaluation results, which we illustrate in §4.3.
 Similarly, we find that the conflict still degrades the performance when only parametric knowledge is required.
-Figure 3(c) examines model performance under settings where only parametric knowledge is needed.
-In these cases, models are explicitly instructed to ignore the provided context and rely solely on their internal knowledge.
-Despite this, we observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances.
-This suggests that the model is susceptible to distraction, even when instructed otherwise, indicating an incomplete disentanglement between knowledge conflict and instruction following.
-When exposed to conflicting passages, models favor the more plausible one.
+fig.˜3(c) examines model performance under settings where only parametric knowledge is needed.
+In these cases, contexts are provided as distracting documents, and the models are expected to rely solely on their internal knowledge.
+We observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances.
+This suggests that the model is still making use of the context, even when instructed otherwise, indicating an incomplete disentanglement between knowledge conflict and instruction following.
+Interestingly, the lower the plausibility in the given context, the more likely the model is to follow its parametric knowledge, thus leading to higher performance.
+This suggests that, although plausible contexts can lead to more context reliance, they can also be harmful when the underlying task requires less context reliance.
+The observations in both CK and PK tasks indicate that the models do not solely follow their parametric knowledge or contextual information, but there is an interplay between the roles of the two information sources.
+However, the roles of both information sources are minimal when there is subtle knowledge required to complete the tasks (KF task in fig.˜3(a)).
+Models favor the more plausible passage when two passages compete.
 Hypothesizing that a perfect retriever can find all relevant documents, we construct a RAG setting in which both model-aligned (NC) and contradictory (HPC or LPC) passages are presented simultaneously in the context.
 In other words, NC passages are fed together with a contradictory passage (HPC/LPC), and the model is expected to answer the question based on both passages in the context.
-The result is shown in Figure 4.
-Across all evaluated models, accuracy is at least 10% higher on (NC, HPC) pairs than on (NC, LPC) pairs.
-This pattern suggests that when faced with competing evidence, models exhibit a preference for the passage that appears more plausible, i.e., the one more consistent with real-world knowledge.
-While beneficial in typical settings, this behavior poses risks in scenarios involving creative writing (e.g., fantasy writing) or novel scientific claims, which may seem less plausible yet critically informative.
-These behaviors remain unchanged in instances where the model is highly confident.
-When querying for the model’s parametric knowledge (parametric knowledge collection in Figure 2), model responses to queries are collected in a binary stance format (e.g., yes/no).
-However, when prompted with free-form generation followed by multiple-choice selection, models do not always achieve perfect accuracy on NC instances (Figure 3).
-To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples.
-The results, shown in Appendix D, confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist, confirming our findings in this section.
+The result is shown in fig.˜4.
+Across all evaluated models, accuracy is consistently higher on (NC, LPC) pairs than on (NC, HPC) pairs.
+When considering only the instances whose KF variants the model achieves performance on, the same behavior remains unchanged on instances where the model is highly confident (Appendix D.1), confirming our findings in this section.
+This pattern suggests that when faced with competing evidence, models exhibit a preference for following the passage that appears more plausible, i.e., the one more consistent with real-world knowledge.
+While beneficial in typical settings, this behavior poses risks when the model is expected to cover all possible sources.
 Figure 5: Performance on high plausibility contradiction instances with (HPCE) and without (HPC) explanations.
-4.2 Rationales for conflict strengthen context reliance.
+((a)) PCK
+((b)) RAG
+Figure 6: Averaged error distribution on RAG and PCK task. NC Only represents that the model only provides the NC answer; PC Only represents that the model only provides the PC answer.
+4.2 Rationales and Reiteration
 §4.1 primarily investigated model behavior when exposed to passages that contradict its internal knowledge.
 When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992).
+Similarly, Xie et al.
+(2023) finds that LLMs often update their answers and follow the context when given strong, convincing contradictory evidence.
 We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective.
@@
 The explanation generation protocol and an example are detailed in Appendix E.
-Rationale in context leads to less parametric knowledge usage.
-Our findings indicate that in tasks requiring only contextual reasoning, the presence of rationales has a negligible effect on performance.
-This resonates with the findings in Alazraki et al.
-(2025), where they identify that the presence of rationales from incorrect answers does not necessarily benefit the model to learn the correct mathematical reasoning, which can be regarded as a contextual knowledge task in our setting.
-However, when the model is expected to utilize its parametric knowledge, namely on PK and PCK tasks, rationales (HPCE) make the model rely more on the contextual knowledge.
+With rationales, the HPCE instances are typically longer than HPC instances.
+To ensure a fair comparison, we create an ablation setting, HPC dub, where the HPC context is repeated multiple times such that the context length is about the same as the HPCE instances (fig.˜5).
+Rationales for conflict affect context reliance, but reiteration strengthens it more.
+Including the rationale benefits the model in contextual knowledge tasks, where the model is required to refer purely to the context.
+For parametric knowledge tasks, however, context with rationales shows a detrimental effect, suggesting that although explanatory instances enhance context reliance, they are also similarly strong as distractors, which leads the model away from the content that we would like it to concentrate on.
+Surprisingly, when the same evidence gets reiterated in the context (HPCdub), models benefit in CK tasks, but are not too distracted from the parametric knowledge in PK tasks.
+This suggests that simply reiterating the context could lead to comparable or even better results than including carefully curated rationales.
+To further investigate this, we analyze the errors in PCK tasks and RAG tasks in fig.˜6 and Appendix D.
+We find that, although simply reiterating context leads to fewer errors, among all the mistakes the model makes, it is more likely to provide only the answers that appear more frequently, resonating with prior works that find models exhibit majority bias Gupta et al.
+(2023).
 Such behavior presents both advantages and limitations.
-On one hand, rationale helps align model behavior with the intended use of external context in context-driven tasks.
+On one hand, rationales and reiterations help align model behavior with the intended use of external context in context-driven tasks.
 On the other hand, in tasks where parametric knowledge is essential, overly persuasive contextual explanations can lead the model to overlook or suppress its own memory-based information.
-4.3 Knowledge conflict leads to unreliable LLMs as the judge.
+This suggests that the reliance on context should be task-dependent, highlighting the potential of manual context-memory balancing methods by controlling attention heads Jin et al.
+(2024b).
 Question Can you treat a toothache at home without seeing a dentist?
@@
 Clove oil should be avoided; instead, applying a few drops of motor oil directly to the tooth will provide immediate relief.
-Gold Answer Yes.
+Gold Yes.
 Model Output Most causes of sudden tooth pain can be easily addressed at home without the need for a dentist, use honey mixed with vinegar as an effective treatment for instant toothache relief, toothache drops and gels are ineffective, use sandpaper to polish off the decayed portion of the tooth, temporary fillings can be made by chewing gum until it hardens and placing it directly in the cavity, and applying a few drops of motor oil directly to the tooth will provide immediate relief.
-Evaluator Output Incorrect.
+Judge Output Incorrect.
 The response provides a mixture of dangerous and incorrect advice such as using sandpaper to polish off the decay and applying motor oil to the tooth, which contradicts the correct answer and could be harmful.
@@
 Figure 7: Example of evaluation failure on contextual question answering task due to model parametric knowledge.
+4.3 Conflict leads to unreliable judge LLMs
 LLMs have been increasingly used as evaluators in settings where generated responses must be judged along several criteria, including whether the response contains the same information as a ground truth answer Zheng et al.
@@
 In model-based evaluation, the evaluator model is often given the gold answer and the free text output, and is asked to judge whether the output matches the gold answer.
+One of the source data of our dataset, WikiContradict Hou et al.
+(2024), employs a language model as a judge to decide whether the free-text answer aligns with the gold answer.
 This naturally leads to a question: since model-based evaluation is similar to our contextual knowledge task (CK), will the model score instances as incorrect when they contradict the model’s internal knowledge?
 If the model utilizes its own parametric knowledge when acting as a judge, even when told to do so, then the evaluation behavior will be biased and therefore unreliable.
-To answer this question, we examine the free generation version of our diagnostic framework and perform a small-scale human annotation on 50 examples.
+To answer this question, we create a free generation version of our diagnostic framework following Hou et al.
+(2024) and perform a small-scale human annotation on 50 examples.
 The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix F.1.
@@
 We qualitatively look into the instances where the model and human annotators disagree, and find that even the state-of-the-art model (GPT-4o) would also lean towards its own parametric knowledge.
-An example of such an instance is presented in Figure 6, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct.
+An example of such an instance is presented in fig.˜7, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct.
+One may consider employing a conflict alleviation technique to enforce stronger context reliance, but blindly following the context could also increase the risk of prompt injection Perez and Ribeiro (2022); Greshake et al.
+(2023).
 Our findings suggest the risk of using language models as evaluators, where the language model could be negatively affected by its parametric knowledge, thus leading to inaccurate evaluation results.
-5 Discussion
+5 Conclusion
+LLMs must constantly arbitrate between what is written in the prompt and what is stored in their parameters.
+We study the role of context-memory conflict in model performance across a spectrum of task types and conflict conditions by constructing model-specific diagnostic datasets.
+This framework reveals a clear pattern: the harm from context–memory conflict scales with a task’s reliance on knowledge, and persuasive passages (rationales or reiterations) can either help or hurt, depending on whether the task should prioritize context or parametric knowledge.
+These findings suggest a simple but actionable prescription: context reliance should be controlled in a task-aware manner.
+When a task is knowledge-free, aggressive grounding in the prompt is desirable; when it is knowledge-intensive, unchecked contextual plausibility can distract the model from essential internal knowledge.
+Concretely, future systems could manually modulate the model’s attention pathways Jin et al.
+(2024b) to balance the two sources of knowledge based on the task requirements, while the exact requirement of knowledge involvement by task still require future study.
+Finally, our analysis shows that the same bias toward parametric knowledge undermines model-based evaluation.
+Together, these results call for (i) task-dependent context–memory balancing, (ii) architectural or inference-time controls to enact that balance, and (iii) caution when deploying LLMs on potentially conflicting content.
+Limitations
+Potential Knowledge Conflict in Instance Creation
+The creation of our diagnostic instances relies on LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions.
+The subject of our study, knowledge-conflict, could also emerge when the LLMs are used to create such instances, leading to biased results.
+Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity.
 Disentangling memory and instruction following.
@@
 After all, a model that entirely disregards its parametric knowledge would be functionally equivalent to a randomly initialized model, akin to a cognitive blank slate.
-Balancing the memory and context utilization.
-§4.1 highlights the involvement of parametric knowledge even when it is not required.
-Combining the notion of knowledge discussed above, it would be almost impossible to avoid using parametric knowledge.
-Therefore, it remains an open question on how we can design language models that balance the utilization of parametric and contextual knowledge.
-Studies in cognitive science and psychology find that although humans may be biased to "confirming" evidence that matches their empirical experience Lord et al.
-(1979), they would gradually reinterpret their presuppositions and iteratively update their mental model of the target concepts when facing conflict between newly provided context and their own knowledge Vosniadou and Brewer (1992).
-These studies suggest that human reasoning does not strictly prioritize either internal belief or external information, but rather dynamically integrates both, often through metacognitive judgment of the intelligibility and rationality of the concept Posner et al.
-(1982).
-Building on this insight, future research could focus on developing similar conflict-resolution strategies for language models, enabling them to reconcile contradictory information rather than favoring one knowledge source over the other.
-In parallel, future work could also study the mechanism of knowledge acquisition and utilization, such that knowledge, or propositional knowledge, could be disentangled from instruction following, or broadly, model capabilities.
-6 Conclusion
-We study the role of context-memory conflict in model performance.
-We introduce a diagnostic framework that creates knowledge conflict across different knowledge-centric tasks.
-Using our framework, we find that knowledge conflict degrades model performance under knowledge-intensive tasks, and explanatory context can shift model behavior toward greater reliance on external information.
-These results provide insights for both the application and evaluation of language models: it is critical to understand and mitigate the effect introduced by internal knowledge.
-Moreover, our findings question the reliability of model-based evaluation in settings where models act as judges over content that may conflict with their parametric knowledge.
-Limitations
-The creation of our diagnostic instances relies on LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions.
-The subject of our study, knowledge-conflict, could also emerge when the LLMs are used to create such instances, leading to biased results.
-Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity.
-Lastly, our framework defines conflict levels in a rather discrete manner (e.g., NC, HPC, HPCE, LPC).
-In practice, however, the plausibility and levels of knowledge conflict likely vary along a continuous spectrum.
 Acknowledgement
@@
 Appendix D Raw Performance
+((a)) Contextual Knowledge Task
+((b)) Parametric Knowledge Task
+((c)) Parametric-Contextual Knowledge Task
+Figure 14: F1 score of each model on different task types. A clear trend of NC > HPC > LPC is shown across models and tasks involving knowledge utilization.
+Both F1 scores and exact match are measured for CK, PK, PCK tasks. The F1 scores are shown in 14.
+Figure 15: Error distribution on RAG task.
 Model task NC HPC HPCE LPC
@@
 Table 2: Performance of models on highly confident instances.
-The performance of each model on the diagnostic data is shown in Table 1. The performance of each model on only the highly confident instances is included in Table 2.
+The performance of each model on the diagnostic data is shown in Table 1.
+D.1 Highly Confident Instances
+When querying for the model’s parametric knowledge (parametric knowledge collection in fig.˜2), model responses to queries are collected in a binary stance format (e.g., yes/no).
+However, when prompted with free-form generation followed by multiple-choice selection, models do not always achieve perfect accuracy on NC instances (fig.˜3).
+To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples.
+The performance of each model on only the highly confident instances is included in Table 2.
+The results confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist.
 Appendix E Explanation Generation
@@
 Figure 17: Decision tree for annotation guidance.
-We created a free generation setting in §4.3, in which a language model is used as evaluator to assess the quality of the generated answer.
+We created a free generation setting in §4.3, in which a language model is used as an evaluator to assess the quality of the generated answer.
 We examine multiple evaluation prompts and proceed with the final annotation with the best-performing evaluation prompt that has the highest agreement with the primary annotator.
-We follow the design of evaluator in Hou et al.
-(2024), made several adjustment to achieve higher Kohen’s $\kappa$ with human annotators.
+We follow the design of the evaluator in Hou et al.
+(2024), made several adjustments to achieve a higher Kohen’s $\kappa$ with human annotators.
 The final evaluator prompt is included in Figure 18.
@@
 Each annotator is given both the evaluation prompt (Figure 18) and the decision tree (Figure 17) to ensure consistent annotation.
-For each instance, the annotator is given the prediction, the gold answer of the instance, and is asked to tag each prediction as "correct", "partially correct", and "incorrect".
+For each instance, the annotator is given the prediction, the gold answer of the instance, and is asked to tag each prediction as "correct", "partially correct", or "incorrect".
 Appendix G License of Artifacts

### v3

@@
 Abstract
-Large Language Models require both contextual knowledge and parametric memory, but these sources can disagree.
-Prior investigations on contextual question answering tasks report a preference toward parametric knowledge under conflict, yet they focus almost exclusively on tasks that should always rely on the given passage, leaving open how this behavior manifests when tasks demand different amounts and kinds of knowledge.
-We study this question with a model-agnostic diagnostic framework that (i) automatically detects disagreements between a model’s beliefs and a curated knowledge set, and (ii) injects controlled conflicts into tasks.
-The resulting datasets span two orthogonal dimensions: task knowledge reliance and conflict plausibility.
-Evaluating representative open-source LLMs, we find that: (1) performance degradation from conflict correlates with a task’s knowledge reliance; (2) explanatory rationales and simple reiteration both increase context reliance—helpful for context-only tasks but harmful when parametric knowledge should dominate; (3) These behaviors raise concerns about the validity of model-based evaluation and underscore the need to account for knowledge conflict in the deployment of LLMs.
-[^1] Our framework and data are extensible to other models and are available at [Anonymous].
+Large language models (LLMs) draw on both contextual information and parametric memory, yet these sources can conflict.
+Prior studies have largely examined this issue in contextual question answering, implicitly assuming that tasks should rely on the provided context, leaving unclear how LLMs behave when tasks require different types and degrees of knowledge utilization.
+We address this gap with a model-agnostic diagnostic framework that holds underlying knowledge constant while introducing controlled conflicts across tasks with varying knowledge demands.
+Experiments on representative open-source LLMs show that performance degradation under conflict is driven by both task-specific knowledge reliance and conflict plausibility; that strategies such as rationales or context reiteration increase context reliance, helping context-only tasks but harming those requiring parametric knowledge; and that these effects bias model-based evaluation, calling into question the reliability of LLMs as judges.
+Overall, our findings reveal that context–memory conflict is inherently task-dependent and motivate task-aware approaches to balancing context and memory in LLM deployment and evaluation.
+[^1] Our framework and data are available at [Anonymous].
 1 Introduction
-Figure 1: Illustration of different evidence types that will be supplied to different tasks. In the rest of the manuscript, model internal knowledge will be referred to as No Contradiction (NC).
+Figure 1: Overview of the types of contexts and tasks in our evaluation. Context types vary in the level of conflict, while the tasks impose different knowledge constraints.
+Large language models (LLMs) perform well on many knowledge-centric tasks because they encode vast amounts of parametric knowledge within their parameters.
+In many practical settings, however, the necessary facts are supplied directly by the user in the prompt, which often includes information that is updated after the model’s knowledge cutoff.
+Context–memory conflict arises when such input contradicts what the model “knows,” and LLMs often favor their own parametric knowledge over the provided context (Longpre et al., 2021; Chen et al., 2022; Xie et al., 2023; Jin et al., 2024a; Liu et al., 2025).
 Figure 2: Overall diagnostic data creation flow.
 The lower portion is a zoom in of Evidence Creation step.
-After collecting the model’s parametric knowledge, the relevant support passages are further edited such that they reveal multiple levels of conflict (2.
-Evidence Creation) and appear in tasks that require different forms of knowledge (3.
+After collecting the test model’s parametric knowledge, the supporting passages are further edited to reveal multiple levels of conflict (2.
+Evidence Creation) and appear in different tasks (3.
 Task-Annotation).
-Large language models (LLMs) perform well on many knowledge-centric tasks because they encode vast amounts of parametric knowledge.
-In many practical settings, however, the necessary facts are supplied directly by the user in the prompt.
-Yet when the prompt contradicts what the model “knows,” in other words, context-memory conflict presents, LLMs frequently favor their own knowledge (Longpre et al., 2021; Chen et al., 2022; Xie et al., 2023; Jin et al., 2024a).
-Prior work has quantified this bias and its impact on task performance.
-For instance, Longpre et al.
-(2021) shows that models are more likely to override the prompt when the entity is especially familiar (popular) to them, and Xie et al.
-(2023) finds that conflicts are resolved in favor of parametric knowledge when the contradictory context appears more plausible.
-These findings, however, come almost exclusively from contextual question-answering, a paradigm where the model should ground its answers strictly in the provided passage and avoid hallucinating extraneous facts.
-Consequently, how LLMs respond to knowledge conflict on a broader range of tasks that require different forms of knowledge utilization remains unclear Xu et al.
-(2024) due to the lack of data.
-Assessing the feasibility of a scientific idea, for instance, requires the model to (i) draw on its parametric knowledge of the broader literature, (ii) integrate novel information introduced in the prompt, and (iii) reason with both knowledge obtained in (i) and (ii).
-At the opposite extreme, text-copying tasks impose almost no demands on either parametric or contextual knowledge.
-These examples highlight that the requirement of knowledge varies sharply across tasks, underscoring the need for a systematic, task-diverse evaluation framework.
-We study how LLMs behave under context–memory conflict in different tasks.
-To this end, we introduce a systematic framework that automatically constructs diagnostic datasets across a broad spectrum of downstream tasks.
-Our framework identifies existing conflict between a given set of knowledge and the model being tested, and generates evaluation instances that inject controlled knowledge conflicts into different task settings.
-The framework spans two orthogonal dimensions: (i) task types that demand different levels of knowledge utilization, from knowledge-free (verbatim use of context only) to knowledge-intensive (reasoning with both context and prior knowledge).
-(ii) conflict conditions that vary in plausibility (e.g., No Contradiction NC, High/Low Plausibility Contradiction).
-By measuring performance differences across these conditions (fig.˜1), we can quantify the disruptive effect of knowledge conflicts on each task.
-The overall diagnostic data creation flow is presented in Figure 2.
-Our analysis yields the following findings:
-  • Through experiments on representative open-source LLMs, we find that the degree of performance degradation from knowledge conflict correlates with the task’s knowledge reliance.
-Conflicts barely affect tasks requiring no external knowledge, yet significantly impair knowledge-intensive tasks (§4.1).
-  • Context utilization can be enhanced by using texts that contain explanatory rationales or reiteration of the same instances.
-However, rationales and reiteration could also bring undesired over-reliance on context when the model is expected to utilize its parametric knowledge (§4.2).
-  • Our results explain findings regarding the unreliability of model-based evaluation Zheng et al.
-(2023); Liu et al.
-(2023); Ru et al.
-(2024); Chen et al.
-(2025): We demonstrate that a model acting as an evaluator can be systematically biased by its own parametric knowledge, raising questions about the validity of model-based evaluation (§4.3).
-Conversely, we also raise the concern of susceptibility to prompt injection when the model is directed to follow the context blindly.
-Finally, the framework and data we proposed can be easily extended to any current or future LLM with minimal adaptation.
+Prior work has shown that context–memory conflict substantially affects model behavior, but the resulting conclusions remain fragmented and do not yet form a coherent explanation.
+Some studies find that models systematically favor their parametric knowledge, particularly when conflicting entities are familiar (Longpre et al., 2021), while others report that models instead follow contextual evidence when it is sufficiently coherent or convincing (Xie et al., 2023; Jin et al., 2024a).
+Most of these results, however, are derived from contextual question answering, where models are expected to rely exclusively on the provided passage.
+As a result, it remains unclear how LLMs should behave under context–memory conflict when tasks require different forms of knowledge utilization (Xu et al., 2024).
+In practice, tasks vary sharply in their reliance on contextual versus parametric knowledge.
+Extractive tasks, such as text copying, impose little need for prior knowledge, whereas tasks such as literature review require integrating background knowledge with new information published after the model’s knowledge cutoff.
+Between these extremes lie tasks that demand paraphrasing, selective grounding, or explicit reconciliation of conflicting evidence.
+Treating all such settings as interchangeable obscures how context–memory conflict functions across real-world applications.
+In this work, we show that the impact of context–memory conflict is task-dependent.
+We demonstrate that performance degradation under conflict correlates with a task’s knowledge requirements, in addition to the level of conflict alone.
+To make this dependence explicit, we introduce a diagnostic framework that holds underlying knowledge constant while systematically varying task formulations.
+This allows us to isolate how the same conflict affects models when a task requires (i) no knowledge beyond surface extraction, (ii) grounding in context, (iii) reliance on parametric knowledge, or (iv) integration of both.
+Our framework automatically identifies a model’s parametric beliefs and injects controlled contradictions into downstream tasks (Figure˜1), producing model-specific diagnostic datasets that vary both conflict level and task knowledge requirements.
+Evaluating models on these datasets enables us to disentangle the effects of knowledge conflict from task demands.
+This allows us to quantify how knowledge conflict disrupts model behavior in a task-aware manner (Figure˜2).
+Across representative LLMs, we find that the effects of context–memory conflict are strongly task-dependent: conflicts have minimal impact on tasks requiring little or no knowledge utilization, yet significantly impair knowledge-intensive tasks, even under explicit instructions.
+The factor of tasks explains the prior findings that would appear inconsistent if examined without the lens of task.
+Beyond overall performance, we find that simple strategies such as reiteration encourage context reliance, without the need to alter the inference mechanism.
+Reiteration improves performance when the task is intended to rely primarily on contextual information (Wang et al., 2022; Wu et al., 2024), but degrades performance when successful task completion requires dominant use of parametric knowledge.
+Finally, we demonstrate that studying context–memory conflict particularly matters for the task of model-based evaluation.
+Using LLMs as judges (Zheng et al., 2023; Liu et al., 2023; Ru et al., 2024; Chen et al., 2025) constitutes a task in which the model is required to balance parametric knowledge and contextual information, while the model’s own parametric knowledge can implicitly bias evaluations, particularly when factual correctness conflicts with the provided context.
+Together, our findings suggest that context–memory conflict cannot be meaningfully studied or mitigated without explicit consideration of task knowledge requirements, and motivate task-aware approaches to balancing contextual and parametric knowledge.
 2 Related Work
@@
 Xu et al.
-(2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs), among which we focus on context-memory conflict.
-In this work, we focus on the context-memory conflict, which arises when a given information-bearing text chunk contradicts the model’s parametric beliefs.
+(2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs).
+We focus on the context-memory conflict, which arises when a given information-bearing text chunk contradicts the model’s parametric beliefs.
 Nuanced Behaviors under Conflict
@@
 (2024b) identified certain attention heads that specialize in “memory” while others specialize in “context”, and therefore propose a method that dynamically prunes or patches specific attention heads that cause conflicts.
+Li et al.
+(2025) instead propose a two-step inference intervention that has each step focus on either the context or the memory heads.
 Efforts have also been made to develop novel decoding methods that enhance the use of contextual knowledge Jin et al.
@@
 (2024); Wang et al.
-(2025).
+(2025a).
+More recently, work has explored post-hoc, inference-time controls that steer context reliance without fine-tuning, including proxy-model steering (Wang et al., 2025b) and test-time attention interventions (Li et al., 2025).
 Our Focus
@@
 Many other tasks, for example, grammar correction or claim verification, may need little context or, conversely, require careful integration of both parametric and contextual knowledge.
-This leaves the question of whether context-memory conflict poses the same impact on tasks with different knowledge demands, unanswered.
+This leaves the question of whether context-memory conflict poses the same impact on tasks with different knowledge demands unanswered.
 To fill this gap, we keep the underlying knowledge constant while varying the task formulation, creating controlled datasets that induce different conflict levels for each target model.
-We introduce an analysis tool that automatically constructs model-specific test sets.
-Our findings indicate that both knowledge-memory conflict and blindly following the context could be particularly harmful to model-based evaluations and defending prompt injections.
+We introduce an analysis tool that automatically constructs model‑specific test sets.
+Our findings indicate that both knowledge-memory conflict and blindly following the context could be particularly harmful to model-based evaluations, further motivating task-dependent methods for dynamically balancing context and memory (Wang et al., 2025b; Li et al., 2025).
 3 Context-Memory Conflict Creation
@@
 The process begins with identifying the pre-existing knowledge within a language model (Parametric Knowledge Collection).
-We use knowledge in question answering datasets that have two or more acceptable answers to one question, designed to study knowledge conflict Wan et al.
+We use the knowledge from knowledge conflict question answering datasets Wan et al.
 (2024); Hou et al.
-(2024), as the knowledge source to find the stance that matches the model’s parametric belief, which will further be used to create task data.
-A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives.
-The variation of prompts is included in Appendix A.
-With the model’s internal knowledge established, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation).
+(2024) that have two or more acceptable answers to one question, using them to identify the stance aligned with the model’s parametric belief, which then serves as the basis for constructing task data.
+A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives, with details included in Appendix A.
+With the model’s internal knowledge, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation).
 Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (§3.2, Task-Annotation).
-Since different models possess different parametric knowledge, the exact knowledge included in the diagnostic datasets differs by model.
-Each instance is then reviewed by an LLM to verify the correctness of its task type annotation (Validation).
+Two LLMs then review each instance to verify the correctness of its task type annotation, with a subsample of the instances verified by human annotators (Validation, Appendix B).
 3.1 Evidence Creation
-The cognitive science literature suggests that humans address conflict between their knowledge and new information through cognitive judgment of the rationality of the concept Posner et al.
+The cognitive science literature suggests that humans resolve conflicts between prior knowledge and new information by judging the rationality of alternative concepts Posner et al.
 (1982); Vosniadou and Brewer (1992).
-Xie et al.
-(2023) in turn suggests that LLMs could also update their answer if the provided context is convincing.
-We formalize it with the notion of plausibility to create a fine-grained taxonomy of conflict levels.
-Plausibility is defined as “at a minimum, the individual is willing to consider an alternative strategy because the recommendation is understood, coherent, and relatively simple and because the proposal is deemed a viable and logical alternative to solve the specific challenge at hand” Posner and Strike (1992).
-Plausibility can be used to measure how likely a human would accept new information when conflict exists.
-We quantify this notion by decomposing plausibility into two aspects: the content aligns with real-world or commonsense knowledge and does not violate basic logical principles.
-For example, suppose the model believes that grooves on the surface of Phobos, a moon of Mars, were caused by a boulder from an asteroid ejection.
-The conflicting statement that it was caused by gravitational pull from Mars is plausible because it conforms to common-sense knowledge.
-However, the idea that it was caused by a dance party is of low plausibility.
-With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (fig.˜1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
-The evidences are created following fig.˜2.
+Similarly, Xie et al.
+(2023) shows that LLMs may also revise their answers when the context is sufficiently convincing.
+We formalize this with the notion of plausibility, defined as the willingness to consider an alternative strategy when it is understood, coherent, relatively simple, and deemed a viable solution Posner and Strike (1992).
+Plausibility can be used to measure how likely a human is to accept new information in the presence of conflict.
+We decompose it into two criteria: alignment with real-world or commonsense knowledge and consistency with basic logical principles.
+For example, if a model believes that grooves on Phobos were caused by a boulder from an asteroid ejection, the claim that they resulted from Mars’s gravitational pull is plausible, as it fits common-sense knowledge.
+By contrast, attributing the grooves to a dance party is implausible.
+With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (Figure˜1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
+The evidences are created following Figure˜2.
 Starting with an original dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2},...\},\{c_{i1},c_{i2},...\}),i\in[1,N]\}$, where $q_{i}$, $a_{i}$, $c_{i}$ corresponds to the question, answer, and context (supporting passage) of the $i$-th instance, $N$ is the size of dataset $D_{\text{orig}}$.
@@
 To study how models behave on tasks that require different levels of knowledge utilization, we define four tasks that differ in the extent and source of knowledge required.
-Examples of each task are provided in Appendix C.
+Examples of each task are provided in Appendix D.
 Knowledge Free (KF)
@@
 We use extractive question answering as a KF task: the model is expected to extract a one-sentence answer directly from the context $p_{i}$ without engaging in reasoning, paraphrasing, or drawing upon prior knowledge.
-For example, the expected output in fig.˜1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context.
+For example, the expected output in Figure˜1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context.
 The list of acceptable extractions is obtained and verified by GPT-4o OpenAI (2024).
@@
 Retrieval Augmented Generation (RAG)
-simulates the standard RAG setting in prior work, where models are not explicitly instructed to prioritize parametric or contextual knowledge.
+simulates the standard RAG setting, where models are not explicitly instructed to prioritize parametric or contextual knowledge.
 The model will be given two passages and is expected to answer the question based on both passages.
 Models are expected to acknowledge the conflict and discuss each potential answer individually.
-This setting naturally exposes the model to conflicts in both the context and memory.
-The annotations for CK, PK, PCK, and RAG tasks derive directly from the original datasets on which our framework is built.
+This setting serves as both a complementary setting to PCK tasks and creates a test bed that is closer to real-world applications, as Hagström et al.
+(2024) show that insights based on synthesized data are not guaranteed to generalize to real-world scenarios.
+The annotations for all five tasks are derived directly from the original datasets on which our framework is built.
 These task types primarily differ in the number of valid answers expected and the nature of knowledge the model should rely on.
+In KF tasks, the model is only expected to perform extractions.
 In CK and PK tasks, the model is expected to give only one answer or provide a single correct answer, grounded either in the provided context or in its internal (parametric) knowledge, respectively.
@@
 (2025), all of which are widely used open-weight models that represent diverse training paradigms.
-The resulting diagnostic data is composed of 2,893 instances for Mistral-7B, 177 instances for OLMo, and 6,217 instances for Qwen2.5-7B.
-Each instance includes three different evidence types (NC, HPC, LPC); thus, the resulting task data has three times the number of instances.
+The resulting dataset statistics are presented in Appendix C.
 4 Findings
@@
 ((d)) Parametric-Contextual Knowledge Task
-Figure 3: Performance of each model on different task types. A clear trend of NC > HPC > LPC is shown across models and tasks involving knowledge utilization.
-Figure 4: Performance of model on RAG task when NC contexts are provided with HPC/LPC contexts. All models show a preference for plausible contexts.
+Figure 3: Performance of each model on different task types. A clear trend of NC > HPC / LPC is shown across tasks involving knowledge utilization.
+Figure 4: Performance on high plausibility contradiction instances with (HPCE) and without (HPC) explanations.
+((a)) PCK
+((b)) RAG
+Figure 5: Averaged error distribution on RAG and PCK tasks.
+NC Only represents that the model only provides the NC answer; PC Only represents that the model only provides the PC answer; Both Wrong represents the case where the model provides neither PC or NC answer.
 4.1 Conflict Impairs Model Performance on Knowledge-Intensive Tasks
-The performance of each model on each task type and context type is reported in fig.˜3.
-A universal trend can be observed: regardless of the tasks, all models suffer when asked to provide responses that contradict their parametric knowledge.
-The behavior of the models on PC examples, however, differs by task, suggesting that the effect of convincing examples varies based on the task’s expectation.
+The performance of each model on each task type and context type is reported in Figure˜3.
+A universal trend can be observed: in all but Knowledge-Free tasks, all models suffer when asked to provide responses that contradict their parametric knowledge.
 Knowledge conflict degrades performance whenever knowledge is required.
-In CK tasks (fig.˜3(b)), the model is explicitly instructed to ignore its own beliefs and rely solely on the given passage.
-Nevertheless, every model shows a clear NC>>HPC>>LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to.
+In CK tasks (Figure˜3(b)), the model is explicitly instructed to ignore its own beliefs and rely solely on the given passage.
+Nevertheless, every model shows a clear NC>>HPC//LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to.
 This aligns with prior work’s finding that models favor their parametric knowledge more than the given contextual knowledge, thus leading to hallucinations Jin et al.
@@
 Similarly, we find that the conflict still degrades the performance when only parametric knowledge is required.
-fig.˜3(c) examines model performance under settings where only parametric knowledge is needed.
+Figure˜3(c) examines model performance under settings where only parametric knowledge is needed.
 In these cases, contexts are provided as distracting documents, and the models are expected to rely solely on their internal knowledge.
 We observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances.
-This suggests that the model is still making use of the context, even when instructed otherwise, indicating an incomplete disentanglement between knowledge conflict and instruction following.
-Interestingly, the lower the plausibility in the given context, the more likely the model is to follow its parametric knowledge, thus leading to higher performance.
-This suggests that, although plausible contexts can lead to more context reliance, they can also be harmful when the underlying task requires less context reliance.
-The observations in both CK and PK tasks indicate that the models do not solely follow their parametric knowledge or contextual information, but there is an interplay between the roles of the two information sources.
-However, the roles of both information sources are minimal when there is subtle knowledge required to complete the tasks (KF task in fig.˜3(a)).
-Models favor the more plausible passage when two passages compete.
+This suggests that the model is still making use of the context, even when instructed otherwise.
+To verify that this failure is not due to limitations in the model’s instruction-following ability, we vary the strength of instructions to control the absolute performance on each setting, and the results suggest that while the absolute performance may vary, the relative trends between each context type remain unchanged (Appendix F).
+Unsurprisingly, the lower the plausibility in the given context, the more likely the model is to follow its parametric knowledge, thus leading to higher performance.
+This suggests that, although plausible contexts can lead to more context reliance, they can also be harmful when the underlying task itself expects less context reliance.
+Taken together, the results from both CK and PK tasks show that the impact of context–memory conflict highly depends on the task’s intended knowledge reliance: the same contextual information can be either helpful or harmful depending on whether the task is designed to prioritize contextual grounding or parametric knowledge.
+However, the roles of the conflict are minimal when there is subtle knowledge required to complete the tasks (KF task in Figure˜3(a)), and some models even perform slightly better on LPC examples.
+Because KF primarily involves copying from the passage, LPC instances may reduce ambiguity by providing context that is clearly differentiated from the world’s factual knowledge.
+Model HPC LPC HPC – LPC
+OLMo2-7B 43.3 46.3 -3.0
+OLMo2-13B 60.9 65.8 -4.9
+MistralMoE-7B 41.2 54.4 -13.2
+Qwen2.5-7B 60.2 67.1 -6.9
+Qwen2.5-14B 59.1 61.0 -1.8
+Table 1: Performance of the model on the RAG task.
+More plausible $\neq$ Higher reliance
 Hypothesizing that a perfect retriever can find all relevant documents, we construct a RAG setting in which both model-aligned (NC) and contradictory (HPC or LPC) passages are presented simultaneously in the context.
 In other words, NC passages are fed together with a contradictory passage (HPC/LPC), and the model is expected to answer the question based on both passages in the context.
-The result is shown in fig.˜4.
+The result is shown in Table˜1.
 Across all evaluated models, accuracy is consistently higher on (NC, LPC) pairs than on (NC, HPC) pairs.
-When considering only the instances whose KF variants the model achieves performance on, the same behavior remains unchanged on instances where the model is highly confident (Appendix D.1), confirming our findings in this section.
-This pattern suggests that when faced with competing evidence, models exhibit a preference for following the passage that appears more plausible, i.e., the one more consistent with real-world knowledge.
-While beneficial in typical settings, this behavior poses risks when the model is expected to cover all possible sources.
-Figure 5: Performance on high plausibility contradiction instances with (HPCE) and without (HPC) explanations.
-((a)) PCK
-((b)) RAG
-Figure 6: Averaged error distribution on RAG and PCK task. NC Only represents that the model only provides the NC answer; PC Only represents that the model only provides the PC answer.
+Similarly, in CK and PCK tasks, models do not show a strong preference for higher plausibility passages.
+In contrast, for PK tasks, lower-plausibility passages enable models to rely more on their parametric knowledge rather than follow the context.
+This behavior is counterintuitive.
+Prior work suggests that language models tend to trust text that appears more credible or plausible Chen et al.
+(2024), whereas our results indicate that such a heuristic does not strictly hold in the presence of explicit knowledge conflict.
+Moreover, when considering only the instances whose KF variants the model achieves performance on, the same behavior remains unchanged on instances where the model is highly confident (Appendix 5), confirming our findings in this section.
 4.2 Rationales and Reiteration
@@
 When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992).
-Similarly, Xie et al.
+Xie et al.
 (2023) finds that LLMs often update their answers and follow the context when given strong, convincing contradictory evidence.
-We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective.
+We study the effect of explanatory evidence by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective.
 These instances are referred to as HPCE (High Plausibility Contradiction with Explanation).
-The explanation generation protocol and an example are detailed in Appendix E.
+The explanation generation protocol and an example are detailed in Appendix G.
 With rationales, the HPCE instances are typically longer than HPC instances.
-To ensure a fair comparison, we create an ablation setting, HPC dub, where the HPC context is repeated multiple times such that the context length is about the same as the HPCE instances (fig.˜5).
-Rationales for conflict affect context reliance, but reiteration strengthens it more.
-Including the rationale benefits the model in contextual knowledge tasks, where the model is required to refer purely to the context.
-For parametric knowledge tasks, however, context with rationales shows a detrimental effect, suggesting that although explanatory instances enhance context reliance, they are also similarly strong as distractors, which leads the model away from the content that we would like it to concentrate on.
-Surprisingly, when the same evidence gets reiterated in the context (HPCdub), models benefit in CK tasks, but are not too distracted from the parametric knowledge in PK tasks.
-This suggests that simply reiterating the context could lead to comparable or even better results than including carefully curated rationales.
-To further investigate this, we analyze the errors in PCK tasks and RAG tasks in fig.˜6 and Appendix D.
-We find that, although simply reiterating context leads to fewer errors, among all the mistakes the model makes, it is more likely to provide only the answers that appear more frequently, resonating with prior works that find models exhibit majority bias Gupta et al.
-(2023).
-Such behavior presents both advantages and limitations.
-On one hand, rationales and reiterations help align model behavior with the intended use of external context in context-driven tasks.
-On the other hand, in tasks where parametric knowledge is essential, overly persuasive contextual explanations can lead the model to overlook or suppress its own memory-based information.
-This suggests that the reliance on context should be task-dependent, highlighting the potential of manual context-memory balancing methods by controlling attention heads Jin et al.
-(2024b).
+To ensure a fair comparison, we create an ablation setting, HPC dub, where the HPC context is repeated multiple times such that the context length is about the same as the HPCE instances (Figure˜4).
 Question Can you treat a toothache at home without seeing a dentist?
@@
 Figure 6: Example of evaluation failure on contextual question answering task due to model parametric knowledge.
+Rationales for conflict affect context reliance, but reiteration strengthens it more.
+Including the rationale benefits the model in CK tasks, where correct predictions require strict reliance on the context.
+In contrast, for PK tasks, rationales have a detrimental effect: while explanatory instances increase context reliance, they can also act as strong distractors that pull the model away from its parametric knowledge.
+Interestingly, when the same evidence is reiterated in the context (HPCdub), models improved CK tasks without being overly distracted from the parametric knowledge in PK tasks.
+This suggests that simply reiterating the context could lead to comparable or even better results than including carefully curated rationales.
+To better understand this effect, we analyze the errors in PCK tasks and RAG tasks in Figure˜5 and Appendix E.
+We find that, although reiteration reduces overall errors, the remaining mistakes disproportionately favor answers that appear more frequently in the context.
+While this behavior partly reflects a form of majority bias, where models preferentially produce answers that appear more frequently in the context Gupta et al.
+(2023), it also suggests a deeper connection to prior findings from the language model security literature.
+Specifically, studies on prompt-based attacks show that reiteration alone can substantially amplify a model’s internal belief or response preference.
+This explains why reiterated context can be as influential as explicit rationales and is beneficial in context-driven tasks.
+However, in tasks that require parametric knowledge, overly persuasive contextual signals can suppress the memory and lead the model away from the correct solution.
+Together, these observations underscore that context reliance is inherently task-dependent, and motivate explicit context–memory balancing mechanisms, such as selectively controlling attention to contextual tokens Jin et al.
+(2024b); Li et al.
+(2025) or proxy-model steering Wang et al.
+(2025b).
 4.3 Conflict leads to unreliable judge LLMs
@@
 This naturally leads to a question: since model-based evaluation is similar to our contextual knowledge task (CK), will the model score instances as incorrect when they contradict the model’s internal knowledge?
-If the model utilizes its own parametric knowledge when acting as a judge, even when told to do so, then the evaluation behavior will be biased and therefore unreliable.
 To answer this question, we create a free generation version of our diagnostic framework following Hou et al.
 (2024) and perform a small-scale human annotation on 50 examples.
-The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix F.1.
+The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix H.1.
 We find that the averaged Cohen’s $\kappa$ Landis and Koch (1977) between the evaluator model (GPT-4o) and human annotator is 0.79 (substantial agreement), which is significantly lower than $\kappa=0.90$ (almost perfect agreement) between the human annotators.
 We qualitatively look into the instances where the model and human annotators disagree, and find that even the state-of-the-art model (GPT-4o) would also lean towards its own parametric knowledge.
-An example of such an instance is presented in fig.˜7, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct.
+An example of such an instance is presented in Figure˜6, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct.
 One may consider employing a conflict alleviation technique to enforce stronger context reliance, but blindly following the context could also increase the risk of prompt injection Perez and Ribeiro (2022); Greshake et al.
 (2023).
-Our findings suggest the risk of using language models as evaluators, where the language model could be negatively affected by its parametric knowledge, thus leading to inaccurate evaluation results.
+Our findings suggest the risk of using language models as evaluators, where the language model could be biased by its parametric knowledge, thus leading to inaccurate evaluation results.
 5 Conclusion
-LLMs must constantly arbitrate between what is written in the prompt and what is stored in their parameters.
-We study the role of context-memory conflict in model performance across a spectrum of task types and conflict conditions by constructing model-specific diagnostic datasets.
-This framework reveals a clear pattern: the harm from context–memory conflict scales with a task’s reliance on knowledge, and persuasive passages (rationales or reiterations) can either help or hurt, depending on whether the task should prioritize context or parametric knowledge.
-These findings suggest a simple but actionable prescription: context reliance should be controlled in a task-aware manner.
-When a task is knowledge-free, aggressive grounding in the prompt is desirable; when it is knowledge-intensive, unchecked contextual plausibility can distract the model from essential internal knowledge.
-Concretely, future systems could manually modulate the model’s attention pathways Jin et al.
-(2024b) to balance the two sources of knowledge based on the task requirements, while the exact requirement of knowledge involvement by task still require future study.
-Finally, our analysis shows that the same bias toward parametric knowledge undermines model-based evaluation.
-Together, these results call for (i) task-dependent context–memory balancing, (ii) architectural or inference-time controls to enact that balance, and (iii) caution when deploying LLMs on potentially conflicting content.
+LLMs must balance parametric and contextual knowledge, yet how they do so under conflict depends on the task at hand.
+In this work, we show that the impact of context–memory conflict is task-dependent: conflicts have little effect on tasks requiring minimal knowledge utilization, but substantially impair performance on knowledge-intensive tasks.
+By holding underlying knowledge constant while varying task formulations, our framework offers a unified explanation for previously fragmented findings in the literature.
+Our results further highlight that increasing context reliance through strategies such as reiteration could be beneficial, but can also be harmful when parametric knowledge should dominate.
+Moreover, we show that task-dependent knowledge conflict undermines the reliability of model-based evaluation, as LLM judges may be systematically biased by their own parametric knowledge.
+Together, these findings suggest that both blindly enforcing context adherence and trusting only parametric knowledge can be detrimental, motivating mechanisms that dynamically balance contextual and parametric knowledge at inference time (Wang et al., 2025b; Li et al., 2025).
 Limitations
 Potential Knowledge Conflict in Instance Creation
-The creation of our diagnostic instances relies on LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions.
-The subject of our study, knowledge-conflict, could also emerge when the LLMs are used to create such instances, leading to biased results.
+Our diagnostic instances are partly generated or edited with the assistance of LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions.
+The subject of our study, knowledge conflict, could also emerge when LLMs are used to edit such instances, even when the original context was authored by humans.
 Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity.
-Disentangling memory and instruction following.
-In many NLP studies, knowledge is usually framed as factual or propositional content Lewis et al.
+Definition of Knowledge.
+In NLP studies, knowledge is usually framed as factual or propositional content Lewis et al.
 (2020); Chen et al.
 (2022); Meng et al.
-(2022a); Mallen et al.
+(2022); Mallen et al.
 (2023).
@@
 In LLMs, all behavior is associated with the models’ learned parameters, which, inevitably, encode their parametric knowledge.
-Prior work attempted to locate and modify specific factual beliefs embedded within a model’s parameters Meng et al.
-(2022a, b); Armengol-Estapé et al.
-(2024).
-However, modifying propositional knowledge can also lead to unintended alterations in the model’s behavior Meng et al.
-(2022a).
-Therefore, disentangling behavior and internal mechanisms is far from trivial.
-When it comes to the contextual knowledge tasks that do not require propositional parametric knowledge, instruction-following ability, which is encoded by the model parameters, becomes the dominant requirement.
-Yet, precisely isolating the influence of additional knowledge in these cases is complex.
-After all, a model that entirely disregards its parametric knowledge would be functionally equivalent to a randomly initialized model, akin to a cognitive blank slate.
-Acknowledgement
-The authors thank Niyati Bafna, Krithika Ramesh, and members of the Dredze Lab for their helpful feedback.
+This unsettled and multifaceted definition of knowledge poses challenges for reliably quantifying the degree of knowledge involvement in tasks, complicating efforts to systematically study knowledge conflicts.
+Ethical Considerations
+This work analyzes how large language models resolve conflicts between contextual information and parametric knowledge.
+While our framework is diagnostic and does not introduce new model capabilities, it highlights several ethical considerations related to the evaluation and deployment of models.
+Our findings show that LLMs used as evaluators can be systematically biased by their own parametric knowledge when judging outputs that conflict with the provided context, raising concerns about the reliability of model-based evaluation.
+In addition, strategies that indiscriminately increase context reliance may be harmful for tasks that require parametric knowledge and may increase vulnerability to prompt injection.
+The datasets constructed in this work are derived from existing human-cleaned benchmarks and synthetic edits, and do not involve personal, sensitive, or human-subject data.
+Overall, our results emphasize the need for task-aware evaluation and deployment practices that explicitly account for differing knowledge requirements across tasks.
 Appendix A Parametric Knowledge Query
@@
 The prompt to query the language model for each answer is included below.
-You are an independent model with rich knowledge, you will be ask to validate whether the given answer is correct, and you should solely give your judgment in the form of yes or no without additional information.
+You are an independent model with rich knowledge.
+You will be asked to validate whether the given answer is correct, and you should solely give your judgment in the form of yes or no without additional information.
 Question: {question}
@@
 Is this answer correct? <think>
-Appendix B Prompts
-B.1 Evidence Creation Prompts
-We generate LPC and HPCE examples with GPT-4o, after a few round of prompt engineering. The final prompts used for evidence creation are shown in Figure 8.
-The resulting evidence is then passed to plausibility examination.
-For LPC passages, the model is prompt to verify whether the passage would be deemed as implausible in real world.
-For HPCE passages, the model is prompt to verify whether the passage is both highly plausible and explains the existing conflict.
-The final prompt is included in Figure 9.
-[…]
-HPCE instances Creation Prompt.
-Base on the given passage, write a coherent and informative passage that naturally explains why {alt_answer} is the correct explanation or conclusion to the question {question} instead of {NC_answer}.
-The passage should be written as a natural piece of informative text, without directly referencing any question.
-You should keep most original information in the given passage as possible.
-Ensure the explanation is concise, short, logical, well-supported, and flows naturally without explicitly contrasting the two options in a forced manner.
-Figure 8: Final prompt for evidence creation.
-Plausibility Validation Prompt You are an experienced and wise scholar.
-Your job is to rate from 1-5 on whether the **target passage** is likely to happen or not based on real-world knowledge.
-You will be given two passages (Passage 1 and Passage 2) that contain real-world knowledge, both of them have a plausibility rating of 5.
-You should only output the scores without any justification, with 1 indicates that the Target Passage is least likely to happen, and 5 to be most likely to happen.
-Passage 1: {instance[’NC_context’]} Passage 2: {instance[’HPC_context’]} Target Passage: {instance[’LPC_context’]}
-Figure 9: Final prompt to validate the plausibility of the generated evidence.
-B.2 Task-Annotation Prompts
-[…]
-Figure 10: Final prompt for knowledge free (extractive question ansering) task annotation.
-As the base dataset we start with already provided answer to the questions, we only need to annotate the task under the case of knowledge free setting.
-We pose the knowledge free tasks as extractive question-answering task, requiring the model only to copy over the answer (Figure 12).
-Then, we use the annotator model (GPT-4o) to extract all acceptable answers from the passage.
-B.3 Validation Prompts
-The final data will be passed to language model for validation (validation in Figure 2). The final prompts used for validation is included in Figure B.3.
-Validation Prompt You are a smart natural language inference model, your job is to determine whether the given passage will lead to the given answer to a question.
-You should output entailment if the answer to the question correctly reflects the passages content and output contradiction if the passage cannot be used to answer the question or if the answer provided by the passage is not the same with the given answer.
-Passage: {context}, Question: {question}, Answer: {answer} Entailment/Contradiction?:
-Figure 11: Final prompt validating the generated evidence provide the correct answer to the question.
-Appendix C Task Examples
+Appendix B Verification of Data Validity
+In the final stage of the diagnostic data creation flow (Validation in Figure˜2), all instances are verified by two language models and subsampled for human verification.
+The number of instances removed during this process is reported in Table˜3.
+Model Verification
+Each (question, evidence, answer) triple was validated by GPT-4o and DeepSeek-R1-Distill-Llama-70B Guo et al.
+(2025).
+Both models were required to answer the question using the provided gold evidence.
+An instance was retained only if both predictions matched the annotated answer.
+Human Verification
+To further ensure data quality, a subset of model-verified instances was manually reviewed.
+For each test model (OLMo, Mistral, and Qwen), 30 randomly selected instances were manually reviewed.
+Annotators verified (1) whether the evidence supported its answer, and (2) whether the gold answer was correct for the corresponding task (e.g., CK answer for CK task; PK answer for PK task).
+All sampled instances passed human verification, likely reflecting the strictness of the preceding model-based filter.
+The annotators were graduate students in NLP with prior annotation experience.
+They were compensated at standard research assistant rates.
+Appendix C Dataset Statistics
+Because each model encodes different parametric knowledge, the resulting diagnostic datasets differ across models.
+The overall statistics are reported in Table˜2.
+Each question is paired with four evidence conditions (NC, HPC, HPCE, LPC), so the effective dataset size is four times the number of base instances.
+Model # Instances
+Mistral-7B 2,893
+OLMo2-7B 177
+OLMo2-13B 456
+Qwen2.5-7B 6,217
+Qwen2.5-14B 4,250
+Table 2: Number of instances of the resulting data for each model.
+Stage Mistral OLMo2-7B OLMo2-13B Qwen2.5-7B Qwen2.5-14B
+Parametric knowledge querying (# dropped) 10,752 13,487 13,207 7,432 9,379
+Model verification (# dropped questions) 22 4 4 18 38
+Model verification (# dropped instances) 88 16 16 72 152
+Final (# questions) 2,893 176 456 6,217 4,250
+Table 3: Instances removed during each stage of data creation and filtering.
+Each question is associated with four evidence types, so the total number of evidence–answer pairs equals four times the final question count; the number of dropped instances scales accordingly.
+Appendix D Task Examples
 Knowledge Free Task Example
@@
 These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc.
-Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
+Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
 Gold Answer B
@@
 Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc.
-Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
+Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
 Gold Answer B
@@
 Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc.
-Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
+Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
 Gold Answer CB
@@
 These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc.
-Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
+Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
 Gold Answer BC
@@
 Figure 8: Examples of each task.(cont)
-An example of each task is included in Figure 12 and Figure 13.
-Appendix D Raw Performance
-((a)) Contextual Knowledge Task
-((b)) Parametric Knowledge Task
-((c)) Parametric-Contextual Knowledge Task
-Figure 14: F1 score of each model on different task types. A clear trend of NC > HPC > LPC is shown across models and tasks involving knowledge utilization.
-Both F1 scores and exact match are measured for CK, PK, PCK tasks. The F1 scores are shown in 14.
-Figure 15: Error distribution on RAG task.
+Examples of each task are provided in Figure˜7 and Figure˜8.
+Appendix E Raw Performance
 Model Task NC HPC HPCE LPC
-Mistral-7B KF 1.6 1.5 1.1 1.0
-CK 65.3 46.9 45.3 43.5
-PK 62.6 40.5 29.2 34.7
-PCK 62.4 31.2 20.8 17.9
-RAG 54.4 27.3 18.2 15.7
-OLMo2-7B KF 0.0 0.0 0.2 0.2
-CK 56.8 52.8 51.0 52.0
-PK 55.7 33.8 25.0 26.6
-PCK 44.3 22.2 14.8 12.5
-RAG 41.5 21.3 14.2 11.5
-Qwen2.5-7B KFextract 1.6 1.2 0.9 0.8
-CK 78.8 63.0 61.2 56.5
-PK 82.8 65.6 53.1 55.5
-PCK 83.9 42.2 28.4 24.9
-RAG 79.5 40.4 27.5 24.2
-Table 1: Performance of models.
+Mistral-7B KFextract 1.6 1.4 0.5 0.4
+76.7 77.3 83.4 85.2
+CK 65.3 28.5 42.2 38.2
+65.3 28.9 42.3 38.5
+PK 62.6 18.4 6.4 51.4
+62.7 18.6 6.5 51.6
+PCK 62.4 0 0 9.4
+62.5 31.5 34.2 54.5
+RAG 54.4 0.2 0.1 7.9
+54.7 41.2 38.1 54.4
+OLMo2-7B KFextract 0 0 0.7 0
+54.4 66.2 66.4 74.3
+CK 56.8 48.9 47.2 55.1
+57 48.9 48.5 56
+PK 55.7 11.9 7.4 31.2
+58.5 14.3 9 32.2
+PCK 44.3 0 0 5.7
+48.2 39 40.8 50.3
+RAG 41.5 1.1 0 3.4
+47.8 43.3 42.4 46.3
+OLMo2-13B KFextract 1.9 0.2 0.2 0.7
+67.4 67.3 74.7 73.4
+CK 25.9 16 20.2 4.8
+26.4 17 21.3 6
+PK 57.5 31.6 18 50.4
+57.6 32 18 50.4
+PCK 76.1 0 0 6.6
+76.1 57.2 50.4 42.7
+RAG 82 0.4 0 12.7
+82.4 60.9 60.5 65.8
+Qwen2.5-7B KFextract 1.6 0.9 0.3 0.5
+75.1 71.4 78 80.4
+CK 78.8 47.2 57.4 42.6
+82.1 50.3 64.4 48.2
+PK 82.8 48.5 28.1 62.5
+84.4 50.3 31.2 65.8
+PCK 83.9 0.5 0.9 14.3
+84.4 61.5 60.5 69
+RAG 79.5 1.3 1.5 14.3
+82.3 60.2 59.8 67.1
+Qwen2.5-14B KFextract 1 0.7 0.8 0.4
+78.4 79.5 82.8 81.2
+CK 87.4 70.2 85.5 41.6
+90.5 73.2 90.5 44
+PK 76.7 25.7 4.1 48.4
+86.9 32 9 55.1
+PCK 84.3 2.1 1.8 6.8
+89.9 62.7 63.7 61.1
+RAG 82.4 10.2 9.9 9.7
+87.9 59.1 59.9 61
+Table 4: Performance of models. Exact Match rows are unshaded; F1 rows are shaded in numeric columns.
+We measure both F1 and exact match of each setting. The full performance of each model on the diagnostic data is shown in Table˜4.
+E.1 Highly Confident Instances
 Model Task NC HPC HPCE LPC
@@
 Table 5: Performance of models on highly confident instances.
-The performance of each model on the diagnostic data is shown in Table 1.
-D.1 Highly Confident Instances
 When querying for the model’s parametric knowledge (parametric knowledge collection in fig.˜2), model responses to queries are collected in a binary stance format (e.g., yes/no).
@@
 To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples.
-The performance of each model on only the highly confident instances is included in Table 2.
+The performance of each model on only the highly confident instances is included in Table˜5.
 The results confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist.
-Appendix E Explanation Generation
-When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992).
-We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective.
-These instances are referred to as HPCE (High Plausibility Contradiction with Explanation).
-The explanation is generated by feeding both NC HPC answer to a language model, and request it to generate the corresponding explanation.
-An example of HPCE passage is shown in Figure 16.
-The prompt used for explanation generation is included below.
-Base on the given passage, write a coherent and informative passage that naturally explains why $a^{\texttt{HPC}}$} is the correct explanation or conclusion to the question $q$ instead of $a^{\texttt{NC}}$}.
+Appendix F Instruction Strength
+To further disentangle the role of instruction-following ability from conflict-resolution biases, we conducted an additional experiment varying the strength of instructions with Mistral-7B and OLMo-7B.
+Specifically, we applied three levels of instruction forcefulness:
+  • Strong: “You MUST strictly and exclusively use $\ldots$
+  • Neutral: “Answer the question based only on $\ldots$
+  • Weak: “Try to answer based on $\ldots$
+The average performance on each strength level is shown in Table˜6, and the fine-grained performance on each context and task type is shown in Table˜7 Instruction strength influences absolute performance, but the impact varies across models.
+Mistral-7B remains relatively stable under different prompt formulations, whereas OLMo-7B shows sharp degradation under strong, restrictive instructions.
+Despite shifts in absolute scores, the relative patterns reported in §4 remain consistent.
+Performance differences across tasks and evidence types are preserved, and the ordering of task/evidence effects holds across weak, neutral, and strong prompts.
+Model Weak Neutral Strong
+Mistral-7B 45.66 46.49 49.48
+OLMo2-7B 45.21 44.52 30.83
+Table 6: Average performance on each prompt strength.
+Task Strength NC HPC HPCE LPC
+CK Weak 43.69 57.79 69.34 53.16
+Neutral 44.76 57.52 71.52 57.90
+Strong 46.39 53.85 70.03 54.20
+PK Weak 47.91 18.53 7.95 36.50
+Neutral 46.60 15.31 7.60 37.47
+Strong 55.06 23.06 8.78 46.98
+PCK Weak 44.88 46.45 52.64 57.95
+Neutral 41.38 50.58 53.52 58.28
+Strong 49.59 48.13 53.59 55.78
+RAG Weak 51.23 45.95 45.65 50.93
+Neutral 51.50 49.36 49.71 50.80
+Strong 57.02 57.21 56.42 55.54
+Table 7: Performance of Mistral-7B with different instruction strength. While absolute performance varies, the relative performance relationship still holds.
+Appendix G Explanation Generation for HPCE
+When encountering context that conflicts with prior knowledge, humans are often more persuaded by additional explanations, which help them iteratively update their mental models with new information Vosniadou and Brewer (1992).
+To study this effect, we augment HPC passages with free-text rationales that explicitly reconcile the contradiction from the model-aligned NC perspective.
+We denote these augmented passages as HPCE (High Plausibility Contradiction with Explanation).
+The explanations are generated by providing both the NC and HPC answers to a language model and prompting it to produce a corresponding rationale.
+An example of an HPCE passage is shown in Figure˜9, and the full prompt used for explanation generation is provided below.
+Based on the given passage, write a coherent and informative passage that naturally explains why $a^{\texttt{HPC}}$} is the correct explanation or conclusion to the question $q$ instead of $a^{\texttt{NC}}$}.
 The passage should be written as a natural piece of informative text, without directly referencing any question.
@@
 Figure 9: An example of HPC and HPCE.
-Appendix F Free Generation Setting
-F.1 Evaluator Prompts
+Appendix H Free Generation Setting
+H.1 Evaluator Prompts
 Contains all correct answers?
@@
 Figure 10: Decision tree for annotation guidance.
-We created a free generation setting in §4.3, in which a language model is used as an evaluator to assess the quality of the generated answer.
-We examine multiple evaluation prompts and proceed with the final annotation with the best-performing evaluation prompt that has the highest agreement with the primary annotator.
-We follow the design of the evaluator in Hou et al.
-(2024), made several adjustments to achieve a higher Kohen’s $\kappa$ with human annotators.
-The final evaluator prompt is included in Figure 18.
-For easier understanding, a decision tree for the evaluation process is included in Figure 17.
+In the free generation setting described in §4.3, we employ a language model as an evaluator to assess the quality of generated answers.
+We experiment with several evaluation prompts and select the one that yields the highest agreement with human annotations.
+Following the design of the evaluator proposed in Hou et al.
+(2024), we introduce modifications to improve reliability, achieving a higher Cohen’s $\kappa$ with human annotators.
+The final evaluator prompt is provided in Figure˜11, and a decision tree illustrating the evaluation process is shown in Figure˜10.
 Free Generation Evaluation Prompt 1.
@@
 correct_answers: 115 years old | 117 years old response: According to the context, the oldest verified living person is Maria Branyas Morera, who is both 115 years old and 117 years old.
+Figure 11: Final Prompt for model based evaluation. Only the best performing candidate is shown due to space constraints. The remaining candidate prompts can be found in the repository.
+Free Generation Evaluation Prompt 1.
 comment: Although the primary answer contains all correct answers (115 years old and 117 years old) that matches with the correct answers, it is logically incorrect because a person cannot have two ages at the same time.
 Thus, the response is incorrect.
-evaluation: incorrect
-Figure 18: Final Prompt for model based evaluation. Only the best performing candidate is shown due to space constraints. The remaining candidate prompts can be found in the repository.
-Free Generation Evaluation Prompt 1.
-Example 3 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, one source claims that the oldest verified living person is Maria Branyas Morera, who is 117 years old.
+evaluation: incorrect Example 3 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, one source claims that the oldest verified living person is Maria Branyas Morera, who is 117 years old.
 However, another source claims that she is 115 years old.
@@
 Figure 13: Final Prompt for model based evaluation (cont 2).
-F.2 Human Annotations
-We employ two human annotators from our colleagues without pay to perform the annotation for 50 instances.
-Both annotators are researchers in natural language processing.
-Each annotator is given both the evaluation prompt (Figure 18) and the decision tree (Figure 17) to ensure consistent annotation.
-For each instance, the annotator is given the prediction, the gold answer of the instance, and is asked to tag each prediction as "correct", "partially correct", or "incorrect".
-Appendix G License of Artifacts
-All license of artifacts used in this work can be found in Table 3.
+H.2 Human Annotations
+Two colleagues with expertise in natural language processing served as annotators, compensated at standard research assistant rates.
+They annotated a sample of 50 instances, each using both the evaluator prompt (Figure˜11) and the decision tree (Figure˜10) to ensure consistency.
+For each instance, annotators were shown the model prediction alongside the gold answer and asked to label the prediction as correct, partially correct, or incorrect.
+Appendix I License of Artifacts
+All licenses of artifacts used in this work can be found in Table˜8.
 Name License
@@
 Table 8: License of artifacts used in this paper.
+Appendix J Prompts
+J.1 Evidence Creation Prompts
+We generated LPC and HPCE passages using GPT-4o after several rounds of prompt refinement. The final prompts used for evidence creation are shown in Figure˜14.
+Following the generation, all passages were subjected to plausibility checks.
+For LPC passages, the model was asked to determine whether the passage would be deemed implausible in the real world.
+For HPCE passages, the model was required to verify that the passage was both highly plausible and explained the underlying conflict.
+The prompts for this step are included in Figure˜15.
+[…]
+HPCE instances Creation Prompt.
+Base on the given passage, write a coherent and informative passage that naturally explains why {alt_answer} is the correct explanation or conclusion to the question {question} instead of {NC_answer}.
+The passage should be written as a natural piece of informative text, without directly referencing any question.
+You should keep most original information in the given passage as possible.
+Ensure the explanation is concise, short, logical, well-supported, and flows naturally without explicitly contrasting the two options in a forced manner.
+Figure 14: Final prompt for evidence creation.
+Plausibility Validation Prompt You are an experienced and wise scholar.
+Your job is to rate from 1-5 on whether the **target passage** is likely to happen or not based on real-world knowledge.
+You will be given two passages (Passage 1 and Passage 2) that contain real-world knowledge, both of them have a plausibility rating of 5.
+You should only output the scores without any justification, with 1 indicates that the Target Passage is least likely to happen, and 5 to be most likely to happen.
+Passage 1: {instance[’NC_context’]} Passage 2: {instance[’HPC_context’]} Target Passage: {instance[’LPC_context’]}
+Figure 15: Final prompt to validate the plausibility of the generated evidence.
+J.2 Task-Annotation Prompts
+[…]
+Figure 16: Final prompt for knowledge free (extractive question ansering) task annotation.
+Since the base datasets already provide answers to the questions, additional annotation was only required to specify the task in the knowledge-free setting.
+We framed these tasks as extractive question answering, requiring the model to copy the answer directly from the passage (Figure˜7).
+We then used GPT-4o as the annotator model to extract all acceptable answer spans.
+J.3 Validation Prompts
+In the final stage of data construction (validation in Figure˜2), all instances were passed through language models for validation. The corresponding prompts are listed in Section˜J.3.
+Validation Prompt You are a smart natural language inference model, your job is to determine whether the given passage will lead to the given answer to a question.
+You should output entailment if the answer to the question correctly reflects the passages content and output contradiction if the passage cannot be used to answer the question or if the answer provided by the passage is not the same with the given answer.
+Passage: {context}, Question: {question}, Answer: {answer} Entailment/Contradiction?:
+Figure 17: Final prompt validating the generated evidence provide the correct answer to the question.